### v3

# A Theoretical Characterization of Optimal Data Augmentations in Self-Supervised Learning

Abstract
Data augmentations play an important role in the recent success of Self-Supervised Learning (SSL). While commonly viewed as encoding invariances into the learned representations, this interpretation overlooks the impact of the pretraining architecture and suggests that SSL would require diverse augmentations which resemble the data to work well. However, these assumptions do not align with empirical evidence, encouraging further theoretical understanding to guide the principled design of augmentations in new domains. To this end, we use kernel theory to derive analytical expressions for data augmentations that achieve desired target representations after pretraining. We consider two popular non-contrastive losses, VICReg and Barlow Twins, and provide an algorithm to construct such augmentations. Our analysis shows that augmentations need not be similar to the data to learn useful representations, nor be diverse, and that the architecture has a significant impact on the optimal augmentations.
Machine Learning, ICML, Theory of Self-Supervised Learning, Kernel Representation Learning, Data Augmentations
1 Introduction
Self-supervised learning (SSL) has gained prominence in recent years, serving as one of the backbones of the foundation models driving current progress in artificial intelligence. Instead of using labels, SSL employs a surrogate objective to learn representations, which are then used for downstream tasks. Joint embedding methods (Chen et al., 2020b; He et al., 2020; Grill et al., 2020; Zbontar et al., 2021) in particular have seen a rise in popularity, achieving competitive performance with supervised representation learning for downstream classification, image segmentation and object detection (Chen & He, 2020; Bardes et al., 2022). Broadly speaking, these models encourage different views of the same underlying point to align closely in the embedding space, while preventing the representations from collapsing (Jing et al., 2022). In the vision domain, commonly used image augmentations are random crop, Gaussian blur, and color distortion (Chen et al., 2020b).
The choice of augmentations is critical to the representations learned by the SSL objective. Different downstream tasks may require different augmentation; for example, Purushwalkam & Gupta (2020) show that cropping encourages invariance to occlusions but negatively affects downstream tasks that require category and viewpoint invariance. In addition, Xiao et al. (2020); Lee et al. (2021); Zhang & Ma (2022); Bendidi et al. (2023) demonstrate that certain augmentations benefit or penalize downstream classification on different classes, and Ericsson et al. (2022) empirically show that pose-related tasks and classification-related tasks benefit from opposite augmentations. Previous theoretical studies largely do not address the subtleties of augmentation choice. Instead, by assuming a certain relationship between the data and the augmentations, they provide guarantees specifically on downstream classification tasks (Arora et al., 2019; HaoChen et al., 2022; Saunshi et al., 2022).
In practice however, the choice of suitable augmentations depends not only on the downstream task, but also intricately on the domain (Bendidi et al., 2023; Balestriero et al., 2023). Medicine still struggles to benefit from the success of SSL, arguably since assumptions that hold for natural images do not hold for medical images (Huang et al., 2023). As an example, consider brain scans. Unlike natural images, which typically include a central object, cropping parts of an image with a tumor can drastically change the interpretation of the scan. In practice, augmentations are therefore carefully crafted for different applications, a manual and empirically driven process. Theoretical insights are scarce, and the few that exist do not explain empirical phenomena well. For example, if augmentations primarily serve the role of encoding invariances, one would expect that augmentations need to (a) capture invariances occuring in the data, and (b) capture a diverse set of such invariances. Neither holds in practice. Firstly, strong data augmentations, such as random crop and cutout, have drastically different marginal distributions than the original data (Gontijo-Lopes et al., 2020). At the same time, these are exactly the augmentations that empirically provide the most benefit for downstream performance (Chen et al., 2020b). In contrast, natural-seeming augmentations, such as adding Gaussian noise, only show limited benefit. Secondly, Bardes et al. (2022) find that it is always better to use more data instead of more augmentations (Cabannes et al., 2023). In fact, Chen et al. (2020b) use only three types of augmentations. Finally, Moutakanni et al. (2024) stretch both of the above assumptions to the limit, showing that with just one augmentation, cropping, one can achieve state-of-the-art representations provided sufficient data.
(a) Previous works
(b) Our study
Figure 1: Previous theoretical works assume certain augmentation characteristics and examine the learned representations. We ask the reverse question: Given target representations (e.g. obtained from a pretrained ResNet), what are the augmentations that achieve these representations?
In self-supervised learning, a unique interplay exists between data, augmentations and representations (Cabannes et al., 2023). Prior works predominantly focus on one side of this interaction, asking how data and augmentations influence the learned representations. In this paper, we study SSL from a different angle. We ask: For given data and desired target representations, what are the augmentations that result in these representations? This conceptually deviates from prior studies, as illustrated in Figure 1.
Contributions.
To tackle this question, we place ourselves within existing theoretical frameworks for SSL. We formally prove that suitable data augmentations together with a sufficiently expressive hypothesis class can learn any desired representation by joint embedding methods such as VICReg (Bardes et al., 2022) and Barlow Twins (Zbontar et al., 2021). We derive analytical expressions for the augmentations in Section 4, and propose an algorithm to compute these augmentations in Section 5. Finally, we give insights about augmentation choice in Section 7, interpreting the aforementioned empirical phenomena through the lens of our analysis. To summarize, our contributions are the following.
  • We prove that for existing theoretical formulations of VICReg and Barlow Twins, it is possible to learn any representation for the input data, given suitable augmentations and a sufficiently expressive function class.
  • For VICReg, we derive a closed-form solution for the augmentations; for Barlow Twins, the augmentations are expressed through a solution of a continuous-time Lyapunov equation. To the best of our knowledge, this is the first method to construct augmentations explicitly for any given target representation.
  • Our theoretical results provide new insights into the role of augmentations in SSL. (a) Augmentations need not capture invariances that occur in the original data. (b) Even very few augmentations can provide good representations, as empirically observed by Moutakanni et al. 2024. (c) Augmentations corresponding to the same representations can be recognizably different depending on the architecture.
2 Related Work
Theory of Self-Supervised Learning.
Previous theoretical works on self-supervised learning primarily focus on mathematically describing the learned representations, while implicitly assuming that useful augmentations are already given. This high-level idea has been formalized in numerous ways. Positive pairs can be assumed to be i.i.d. samples from the same class (Arora et al., 2019), or a data augmentation graph is defined in which positive pairs are connected (HaoChen et al., 2022). Furthermore, a suitable integral operator can encode similarity between positive pairs (Cabannes et al., 2023), or the target representations can assumed to be contained in a Hilbert space defined by the augmentation (Zhai et al., 2024). In this work, we turn this question around: Given target representations, what are the augmentations one needs to learn these representations?
Kernel Methods and SSL.
Studying supervised deep learning through kernel methods has proven to be fruitful, most notably by virtue of the neural tangent kernel (NTK) (Jacot et al., 2018). This has sparked several works that assume kernel models for the representation function in SSL (Kiani et al., 2022; Balestriero & LeCun, 2022; Simon et al., 2023; Cabannes et al., 2023; Esser et al., 2024). In this paper, we also consider a kernel setting. Our framework for VICReg is equivalent to that of Balestriero & LeCun (2022), while for Barlow Twins, our framework is similar to that of Simon et al. (2023).
The Role of Data Augmentations.
The role of data augmentations has been studied both in the supervised setting as well as in the self-supervised setting. In the supervised setting, Gontijo-Lopes et al. (2020); Kim et al. (2022) suggest that augmentations should be similar to the original data distribution, and diverse. Chen et al. (2020a) view augmentations as group actions that keep the data distribution approximately similar. Geiping et al. (2023) show that augmentations can be helpful even if they encode invariances not present in the original data. In SSL, Wang & Isola (2020) explain the role of data augmentations through the alignment of the positive samples and the uniformity of induced distribution. Wang et al. (2022) present the role of augmentations as connecting the data through overlaps, which then cause class-separated representations in downstream classification. Kügelgen et al. (2021) view augmentations as transformations that preserve the semantic content of an image while changing the style component. Taking an information theoretic perspective, Tian et al. (2020) argue that augmentations should reduce the mutual information between views as much as possible while keeping the shared task-relevant information intact. It is challenging (and in some cases even impossible) to reconcile these interpretations with the empirical observation that augmentations produce useful representations even when they are not diverse or similar to the original distribution (Moutakanni et al., 2024).
3 Preliminaries
3.1 Problem Statement and Approach
Joint Embedding Loss Functions.
Throughout this paper, we assume we are given $n$ data points $\{x_{i}\}_{i=1}^{n}\subseteq\mathcal{X}\subseteq\mathbb{R}^{m}$. For each of these points $x_{i}$, we create two views by sampling from a random augmentation map $T:\mathcal{X}\rightarrow\mathcal{X}$ with distribution $\mathcal{T}$. This yields $T(x_{i})$ and $T^{\prime}(x_{i})$. Both views are passed to a function $f:\mathcal{X}\rightarrow\mathbb{R}^{d}$ that maps to a lower-dimensional space with $d<\min(n,m)$. In practice, $f$ is typically a neural network, or a kernel function. We denote $\mathcal{F}$ for the function class from which $f$ is chosen. Denoting $z_{i}=f(T(x_{i}))$ and $z_{i}^{\prime}=f(T^{\prime}(x_{i}))$, a joint embedding loss $L(Z,Z^{\prime})$ is computed on the matrices $Z=[z_{1},\ldots,z_{n}]$ and $Z^{\prime}=[z_{1}^{\prime},\ldots,z_{n}^{\prime}]\in\mathbb{R}^{d\times n}$. Possible choices of $L$ will be discussed later. The loss to be minimized by the model $f$ is
($\mathcal{L}\Bigl{(}\{x_{i}\}_{i=1}^{n},\mathcal{T},f\Bigl{)}=\mathbb{E}_{Z,Z^{\prime}}\Bigl{[}L(Z,Z^{\prime})\Bigl{]}$
where the expectation is with respect to the randomness in the augmentations. Depending on the method, $T$ and $T^{\prime}$ can either be independent samples from the underlying distribution $\mathcal{T}$, or they can be conditioned on the event $T^{\prime}\neq T$ (in that case, augmentations always yield two distinct views of the same point $x_{i}$).
RKHS and Function class.
We assume that we are given a kernel $\kappa:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. The canonical feature map of this kernel $\kappa$ is denoted by $\phi(x):\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is the reproducing kernel Hilbert space (RKHS) associated with $\kappa$. From there, we adopt the notation of Cabannes et al. (2023) and use the tensor product notation $\mathbb{R}^{d}\otimes\mathcal{H}$ to denote the set of linear maps from $\mathcal{H}$ to $\mathbb{R}^{d}$. An element $\Theta\in\mathbb{R}^{d}\otimes\mathcal{H}$ can be thought of as a matrix with $d$ rows that are elements of $\mathcal{H}$. The adjoint operator, denoted as $\Theta^{T}\in\mathcal{H}\otimes\mathbb{R}^{d}$, can be thought of as the transposed matrix. We equip the function space with the Hilbert Schmidt norm on $\mathbb{R}^{d}\otimes\mathcal{H}$. The space $\mathcal{F}$ from which our model $f$ is chosen is given by the set of all functions $x\mapsto\Theta^{T}\phi(x)$ where $\Theta\in\mathbb{R}^{d}\otimes\mathcal{H}$, again equipped with the Hilbert Schmidt norm on $\Theta$. In this paper, we restrict ourselves to the least norm solutions of ($\mathcal{L}\left(\{\phi(x_{i})\}_{i=1}^{n},\mathcal{T},f\right)$ on $\mathcal{F}$, as is commonly done in kernel methods. Formally:
Definition 3.1 (Least Norm Minimizers).
Let $\mathcal{L}$ be a joint embedding loss. A function $\hat{f}$ is said to minimize $\mathcal{L}$ with least norm over $\mathcal{F}$ if $\mathcal{L}(\hat{f})=\inf_{f\in\mathcal{F}}\mathcal{L}(f)$ and $\|\hat{f}\|_{\mathcal{F}}\leq\|f\|_{\mathcal{F}}$ for all $f$ with $\mathcal{L}(f)=\inf_{g\in\mathcal{F}}\mathcal{L}(g)$.
Problem Statement.
Given data $\{x_{i}\}_{i=1}^{n}\subset\mathcal{X}$ and a desired target function $f^{\ast}:\mathcal{X}\rightarrow\mathbb{R}^{d}$, our goal is to find a distribution $\mathcal{T}$ such that $f^{*}$ is equivalent to $\min_{f\in\mathcal{F}}\mathcal{L}(\{x_{i}\}_{i=1}^{n},\mathcal{T},f)$. We consider two functions $f$ and $g$ to be equivalent if they are identical up to an invertible affine transformation. This notion of equivalence is justified by HaoChen et al. (2022, Lemma 3.1), who prove that invertible affine transformations do not influence the downstream performance when using the standard linear evaluation protocol (Chen et al., 2020b).
Figure 2: An illustration of our approach. We find a transformation $T_{\mathcal{H}}$ in the RKHS that yields the target representations $f^{*}$ (see Theorems 4.2 and 4.4). We then translate $T$ to the input space by solving a pre-image problem $\tilde{\phi}^{-1}$ (Algorithm 1).
Our Approach.
Motivated by the empirical study of Moutakanni et al. (2024), we consider distributions $\mathcal{T}$ that can lead to only two possible augmentations: The identity and some other transformation $T$. We construct $T$ by first finding a suitable augmentation in the Hilbert space $\mathcal{H}$, and only later translating it back to the input space $\mathcal{X}$. Given a target representation $f^{*}$, we first identify a binary distribution $\mathcal{T}_{\mathcal{H}}$ that either samples from the identity operator in $\mathcal{H}$ or reduces to a transformation $T_{\mathcal{H}}$. The distribution $\mathcal{T}_{\mathcal{H}}$ is carefully chosen in a way that ensures the least-norm minimizer of the loss is equivalent to the desired target $f^{*}$. Having found the correct augmentation in the Hilbert space, we translate it back to the input space by defining $T\coloneqq\tilde{\phi}^{-1}\circ T_{\mathcal{H}}\circ\phi$ where $\tilde{\phi}^{-1}$ is a solution to the pre-image problem for kernel machines. Our approach is illustrated in Figure 2.
Continuous-Time Lyapunov Equations.
Our theoretical results for the Barlow Twins loss partly build on solutions to continuous-time Lyapunov equations. These are matrix equations of the form $AX+XA^{T}=C$, where $X,A,C$ are matrices of appropriate size. For our purposes, the main result we need is that if $C$ is symmetric and $A$ is a positive-definite matrix, then there is a unique symmetric solution $X$ that solves the Lyapunov equation (See Ortega, 1987, Theorem 6.4.2). The solution $X$ can be represented analytically in terms of $C$ and $A$.
Notation.
For points $\{x_{i}\}_{i=1}^{n}$ we denote by $\Phi\in\mathcal{H}\otimes\mathbb{R}^{n}$ the operator $[\phi(x_{1}),\ldots,\phi(x_{n})]$. $\Phi$ can be thought of as the transpose of the design matrix. We use $\mathcal{D}\otimes\mathcal{D}$ to denote the product of two probability measures $\mathcal{D}$. $I_{\mathcal{H}}$ is the identity operator. $\circ$ is the composition operator. $I_{n}\in\mathbb{R}^{n\times n}$ denotes the identity matrix, and $\mathbf{1}_{n}\in\mathbb{R}^{n}$ denotes a vector of ones. The centering matrix is defined as $H_{n}\coloneqq I_{n}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$. We omit subscripts where the dimension is clear from the context. We denote the covariance matrix by $\text{cov}(X)=\frac{1}{n}(XH)(XH)^{T}$. Finally, $O(n)$ denotes the set of $n\times n$ orthonormal matrices.
3.2 Assumptions
In this paper, we assume two conditions to be satisfied. Firstly, the target representations should have full rank. Secondly, the Gram matrix of the kernel should be invertible. For universal kernels (e.g. the Gaussian RBF kernel), Condition 3.3 holds for any set of distinct samples $\{x_{i}\}_{i=1}^{n}$.
Condition 3.2.
Given a target function $f^{*}$ and data $\{x_{i}\}_{i=1}^{n}$, the covariance matrix $\text{cov}\left([f^{\ast}(x_{1}),\ldots,f^{\ast}(x_{n})]\right)$ has full rank $d$.
Condition 3.3.
Given a kernel function $\kappa(\cdot,\cdot)$ and data $\{x_{i}\}_{i=1}^{n}$, the Gram Matrix $K=[\kappa(x_{i},x_{j})]_{i,j}$ has full rank.
3.3 Non-Contrastive Loss Functions
VICReg.
The VICReg loss (Bardes et al., 2022) is defined via three components.
$s(Z,Z^{\prime})=$ $\frac{1}{n}\sum_{i=1}^{n}\left\|z_{i}-z_{i}^{\prime}\right\|_{2}^{2}$
$v(Z)=$ $\frac{1}{d}\sum_{i=1}^{d}\left(1-[\text{cov}\left(Z\right)]_{i,i}\right)^{2}$
$c(Z)=$ $\frac{1}{d}\sum_{i\neq j}^{d}[\text{cov}(Z)]_{i,j}^{2}$
[…]
$L_{\text{VIC}}(Z,Z^{\prime})=$ $\lambda s(Z,Z^{\prime})+\mu\Bigl{[}v(Z)+v(Z^{\prime})\Bigl{]}$
$+\nu\Bigl{[}c(Z)+c(Z^{\prime})\Bigl{]}$
Finally, the VICReg loss is defined as
$\mathcal{L}_{\text{VIC}}\Bigl{(}\{x_{i}\}_{i=1}^{n}\mathcal{T},f\Bigl{)}=\mathbb{E}_{Z,Z^{\prime}}\Bigl{[}L_{\text{VIC}}(Z,Z^{\prime})\Bigl{]}$
$z_{i}=f\Bigl{(}T(x_{i})\Bigl{)},z_{i}^{\prime}=f\Bigl{(}T^{\prime}(x_{i})\Bigl{)}$, where $\left(T,T^{\prime}\right){\sim}\mathcal{T}\otimes\mathcal{T}$.
This definition is identical to the original VICReg loss (Zbontar et al., 2021), except we use the $L_{2}$ loss instead of the hinge loss in the definition of $v$. To be consistent with previous theoretical works (Balestriero & LeCun, 2022; Cabannes et al., 2023), we also define $v$ using the variance instead of the standard deviation, however our results still hold when using the standard deviation (Appendix A).
Barlow Twins.
The Barlow Twins Loss (Zbontar et al., 2021) is defined in the following way:
$L_{\text{BT}}(Z,Z^{\prime})$ $=\sum_{i}^{n}(1-\mathcal{C}_{ii})^{2}+\lambda\sum_{i\neq j}(1-\mathcal{C}_{ij})^{2}$
Where $\lambda>0$ is a hyper-parameter, and the matrix $\mathcal{C}$ depends on the views of the data. Similar to Simon et al. (2023), we consider a symmetrized version of the unnormalized cross-correlation:
$\mathcal{C}=\frac{1}{2n}(ZZ^{\prime T}+Z^{\prime}Z^{T})$
The overall loss is:
$\mathcal{L}_{\text{BT}}\Bigl{(}\{x_{i}\}_{i=1}^{n},\mathcal{T},f\Bigl{)}=\mathbb{E}_{Z,Z^{\prime}}\Bigl{[}L_{\text{BT}}(Z,Z^{\prime})\Bigl{]}$
$z_{i}=f\Bigl{(}T(x_{i})\Bigl{)},z_{i}^{\prime}=f\Bigl{(}T^{\prime}(x_{i})\Bigl{)}$, where $\left(T,T^{\prime}\right){\sim}\mathcal{T}\otimes\mathcal{T}$ conditioned on the event $T\neq T^{\prime}$.
4 Main Results
In this section, we develop our main results for VICReg (Theorem 4.2) and Barlow Twins (Theorem 4.4), proving that any desired target representations can be found by using suitable augmentations, up to equivalence. All proofs are included in Appendix B. We begin by defining the optimal augmentation for VICReg, and then verify its optimality.
Definition 4.1 (Optimal VICReg Augmentations).
Consider a target representation of the form $f^{*}=C\Phi^{T}$, where $C\in\mathbb{R}^{d\times n}$ has full rank. Assuming Condition 3.3, we define $\mathcal{T}_{\mathcal{H}}^{\text{VIC}}(C)$ as a distribution of transformations that applies $I_{\mathcal{H}}$ and $T_{\mathcal{H}}^{\text{VIC}}(C)$ with probability $\frac{1}{2}$ each, where
$T_{\mathcal{H}}^{\text{VIC}}(C)\;:=\;\Phi\,C^{T}\bigl{(}C\,K\,C^{T}\bigr{)}^{-1}C\,\Phi^{T}.$
Theorem 4.2 (Optimality of Augmentations for VICReg).
Let $f^{*}=C\,\Phi^{T},C\in\mathbb{R}^{d\times n}$ satisfy Condition 3.2. Assume Condition 3.3. Then, $C$ has full rank and for any $f$ that is a least norm minimizer of
$\mathcal{L}_{\text{VIC}}\Bigl{(}\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}^{\text{VIC}}(C),\;f\Bigr{)}$
over $\mathbb{R}^{d}\otimes\mathcal{H}$, there exist an invertible matrix $M\in\mathbb{R}^{d\times d}$ and a vector $b\in\mathbb{R}^{d}$ such that
$f\;=\;M\,f^{*}\;+\;b.$
Our main results for Barlow Twins take a similar form. We again state the optimal augmentations first.
Definition 4.3 (Optimal Barlow Twins Augmentations).
Given a rank $d$ matrix $C\in\mathbb{R}^{d\times n}$ and assuming Condition 3.3, we define $\mathcal{T}_{\mathcal{H}}^{\text{BT}}(C)$ as a distribution of transformations that applies $I_{\mathcal{H}}$ and $T_{\mathcal{H}}^{\text{BT}}(C)$ with probability $\frac{1}{2}$ each, where:
$T_{\mathcal{H}}^{\text{BT}}(C)\;:=\;\Phi K^{-\frac{1}{2}}BK^{-\frac{1}{2}}\Phi^{T}.$
Here, $B$ is the unique solution to the continuous-time Lyapunov equation
$KB+BK^{T}=2n\cdot K^{\frac{1}{2}}C^{T}\Bigl{(}CKC^{T}\Bigr{)}^{-2}CK^{\frac{1}{2}}.$
Theorem 4.4 (Optimality of Augmentations for Barlow Twins).
Let $f^{*}=C\,\Phi^{T},C\in\mathbb{R}^{d\times n}$ satisfy Condition 3.2. Assume Condition 3.3. Then, $C$ has full rank and for any $f$ that is a least norm minimizer of
$\mathcal{L}_{\text{BT}}\Bigl{(}\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}^{\text{BT}}(C),\;f\Bigr{)}$
over $\mathbb{R}^{d}\otimes\mathcal{H}$, there exist an orthogonal matrix $Q\in O(d)$ such that
$f\;=\;Qf^{*}.$
Remark.
Assuming the functional form $f^{*}=C\Phi^{T}$ for the target representation is not a restrictive condition. By virtue of the representer theorem (Schölkopf et al., 2001), any least norm minimizer of a loss function that only depends on the training data is certainly contained in the span of $\{\phi(x_{i})\}_{i=1}^{n}$. For a general function $f^{*}:~{}\mathcal{X}\rightarrow\mathbb{R}^{d}$, for example a pretrained ResNet, Proposition 4.5 below allows exactly reconstructing representations $\{f^{*}(x_{i})\}_{i=1}^{n}$ on any set of training data.
Proposition 4.5 (Reconstruction of General Representations).
Let $(\mathcal{L},\mathcal{T_{H}})\in\Bigl{\{}(\mathcal{L}_{\text{VIC}},\mathcal{T}_{\mathcal{H}}^{\text{VIC}}),(\mathcal{L}_{\text{BT}},\mathcal{T}_{\mathcal{H}}^{\text{BT}})\Bigr{\}}$ and $f^{*}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ be a function that satisfies Condition 3.2 on data $\{x_{i}\}_{i=1}^{n}$ satisfying Condition 3.3. Define $F=\Bigl{[}f^{*}(x_{1}),\ldots,f^{*}(x_{n})\Bigr{]}$. Then, for any $f$ that is a least norm minimizer of
($\mathcal{L}\Bigl{(}\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(FK^{-1}),\;f\Bigr{)}$
over $\mathbb{R}^{d}\otimes\mathcal{H}$, there exist an invertible matrix $M\in\mathbb{R}^{d\times d}$ and a vector $b\in\mathbb{R}^{d}$ such that
$\forall_{i\in[n]}{(f\circ\phi)(x_{i})=Mf^{\ast}(x_{i})+b}$
5 Algorithm
In the previous section, we derived $T_{\mathcal{H}}$ as maps in the RKHS $\mathcal{H}$, taking the form of $\phi(x_{i})\mapsto\Phi M\Phi^{T}\phi(x_{i})$ for a matrix $M\in\mathbb{R}^{n\times n}$. In particular, this construction yields augmented versions of original points $\phi(x)$ that are of the form $\Phi\theta$ for some $\theta\in\mathbb{R}^{n}$, and as such lie in the Hilbert space. In this section, we translate these augmentations back to the input space by identifying $x^{\prime}\in\mathcal{X}$ such that $\phi(x^{\prime})\approx\Phi\theta$. In general, such $x^{\prime}$ need not exist — a well known fact that is known as the pre-image problem for kernel machines.
The Pre-Image Problem in Kernel Machines.
We need to find $x^{\prime}\in\mathbb{R}^{m}$ such that $\phi(x^{\prime})$ approximates $\Phi\theta$. An exact solution is not guaranteed to exist (Mika et al., 1998) but can be approximated (Kwok & Tsang, 2004). For our purposes, we use the closed-form approximation proposed by Honeine & Richard (2011): Given the training data in the input space $X=[x_{1},\ldots,x_{n}]\in\mathbb{R}^{m\times n}$ and a single point $\phi^{\prime}=\Phi\theta$ in the Hilbert space, we solve
$\tilde{\phi}^{-1}(\theta)=\arg\min_{x^{\prime}}\Bigl{\|}X^{T}x^{\prime}-(X^{T}X-\mu_{P}K^{-1})\theta\Bigr{\|}^{2}$
where $\mu_{P}>0$ is a hyperparameter and $K$ is, as usual, the kernel Gram matrix on $X$. We also emphasize that while our augmentations $T_{\mathcal{H}}$ are formally defined in the Hilbert space (and may be of independent interest as objects in themselves), the kernel trick allows all computations to happen on matrices in the input space. Moreover, for new data points $x^{\prime}$, the functional form of the augmentations allows evaluating the transformation of $\phi(x^{\prime})$.
Kernel Ridge Regression.
As mentioned in the previous section, our method allows exactly matching any desired target representation $f^{*}$ on given training data $\{x_{i}\}_{i=1}^{n}$ up to equivalence (i.e. up to an invertible affine transformation). It is natural to question the statistical soundness of interpolating on target representations $f^{*}(x_{i})$ obtained e.g. from a ResNet. Therefore, we also incorporate the possibility of using a ridge parameter $\lambda>0$ to avoid overfitting. We solve
$\arg\min_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\left\|f^{*}(x_{i})-f(x_{i})\right\|^{2}+\lambda\|f\|_{\mathcal{F}}^{2}$
where $\mathcal{F}$ is, as before, the space of functions of the form $x\mapsto\Theta^{T}\phi(x),\Theta\in\mathbb{R}^{d}\otimes\mathcal{H}$ equipped with the Hilbert Schmidt norm on $\Theta$. The optimal solution is
$f=$ $F(K+\lambda_{\text{ridge}}I)^{-1}\Phi^{T}$
$F\coloneqq$ $\Bigl{[}f^{*}(x_{1}),\ldots,f^{*}(x_{n})\Bigr{]}$
[…]
Scalability.
Much work has been done to make kernel methods scalable. Of particular importance are random feature methods (Rahimi & Recht, 2007) and the Nyström method (Williams & Seeger, 2000). Kernel ridge regression has been adapted to utilize GPU hardware (Meanti et al., 2020, 2022). These methods propose approximators of the form $f=C[\phi(x_{1}),\ldots,\phi(x_{n^{\prime}})]^{T}$ where where $C$ is calculated as an efficient solution to KRR for $n$ points, $n^{\prime}<n$. These approximate solutions are still of the functional form necessitated by Theorems 4.2 and 4.4, and so we can use them to scale our approach to large datasets.
Algorithm 1 Augmentation Learning for SSL
[…]
0:  Augmented data $\{\hat{x}^{\prime}_{i}\}_{i=1}^{k}$
1:  $F\leftarrow[f^{*}(x_{1}),\ldots,f^{*}(x_{n})]$
2:  $K\leftarrow$ Kernel matrix between $\{x_{i}\}_{i=1}^{n}$ and $\{x_{i}\}_{i=1}^{n}$.
3:  $K_{X,\hat{X}}\leftarrow$Kernel matrix between $\{x_{i}\}_{i=1}^{n}$ and $\{\hat{x}_{i}\}_{i=1}^{k}$.
4:  Solve KRR to obtain representer coefficients:
$C\leftarrow F(K+\lambda_{\text{ridge}}I)^{-1}$
5:  if method is VICReg then
6:     Compute augmentation coefficients (Thm. 4.2):
$C_{aug}\leftarrow C^{T}\left(CKC^{T}\right)^{-1}CK_{X,\hat{X}}$
7:  else if method is Barlow Twins then
8:     $B\leftarrow$ Solution of the Lyapunov equation:
$KB+BK^{T}=2nK^{\frac{1}{2}}C^{T}(CKC^{T})^{-2}CK^{\frac{1}{2}}$
9:     Compute augmentation coefficients (Thm. 4.4):
$C_{aug}\leftarrow K^{-\frac{1}{2}}BK^{-\frac{1}{2}}K_{X,\hat{X}}$
10:  end if
11:  return solution of the pre-image problem with $C_{aug}$.
6 Experiments
We apply Algorithm 1 to MNIST (LeCun et al., 2010) using representations obtained from a ResNet50 architecture pretrained on ImageNet. We show the results for VICReg using different kernel functions in Figure 3. It can be observed that the marginal distribution of the augmented images is different from the original data, and that different kernel architectures result in semantically different augmentations. This is a direct consequence of the analytical form of the augmentations, which depends on the kernel matrix $K$.
To further verify our theoretical results, we measure the similarity between the target function $f^{*}$ and a minimizer $f$ of $\mathcal{L}$, under the augmentation $\mathcal{T_{H}}$. The proof of Theorem 4.2 shows that the affine transformation between $f$ and $f^{*}$ takes the form $f=WQf^{\ast}+b$, where $W$ is a whitening matrix, $b$ is the mean of the representations and $Q\in O(d)$ is an orthogonal matrix. Thus, in the case where $\text{cov}([f^{*}(x_{1}),\ldots,f^{*}(x_{n})])=I_{d}$ and the representations are centered, we would expect $f=Qf^{*}$. We verify this in our next set of experiments by computing the average Procrustes distance between $\{f^{*}(x_{i})\}_{i=1}^{n}$ and $\{f(x_{i})\}_{i=1}^{n}$. It is defined as:
$\min_{Q\in O(d)}\frac{1}{n}\|F-QF^{*}\|_{F}$
where $\|\cdot\|_{F}$ is the Frobenius norm, $F^{*}=[f^{*}(x_{1}),\ldots,f^{*}(x_{n})]$ and $F=[f^{*}(x_{1}),\ldots,f^{*}(x_{n})]$.
Figure 3: Comparison of original and augmented MNIST images for different function classes (defined by the kernel). The target representations are given by the last layer of a ResNet50 pretrained on ImageNet. Notably, different function classes require different augmentations to achieve the same representations.
(a) Theoretical VICReg
(b) VICReg (Bardes et al., 2022)
Figure 4: The average Procrustes distance between the learned representations and target / random representations during training given the augmentations $\mathcal{T}_{\mathcal{H}}^{\text{VIC}}(FK^{-1})$. We consider both the VICReg loss as defined in Section 3 and the original VICReg loss. The augmentations achieve the target representations up to equivalence for both losses.
In Figure 4, we optimize $\mathcal{L}_{\text{VIC}}$ under the augmentation $\mathcal{T}_{\mathcal{H}}^{\text{VIC}}(F^{\ast}K^{-1})$ with target representations obtained by using a ResNet50 (He et al., 2016) pretrained on ImageNet (Deng et al., 2009). We compute the target representations for two classes of MNIST (LeCun et al., 2010) and use the RBF kernel. We compare the Procrustes distance of the learned representations to the target representations, and of the learned representations to random representations with the same covariance. We also run this experiment using the original VICReg loss (Bardes et al., 2022). We observe that the representations learned with the constructed augmentation indeed achieve the target representations. The final average Procrustes distance to the target representations after 3,000 epochs is 0.4 ± 0.1 for the theoretical VICReg loss and 0.5 ± 0.1 for the original VICReg loss. Details on the experiments throughout this section can be found in Appendix C.
7 Discussion
The Role of Augmentations.
Augmentations are commonly thought of as encoding invariances into the representations learned in SSL. Our theoretical analysis provides a more nuanced look on this matter. Theorems 4.2 and 4.4 show that an appropriate choice of augmentation “guides” the solution to a specific low-dimensional subspace of $\mathbb{R}^{k}\otimes\mathcal{H}$. Moreover, both Theorems 4.2 and 4.4 suggest that even single augmentations can learn good representations, and this augmentation need not align with the original data distribution. In fact, for VICReg the augmentation in Definition 4.1 has a provably different marginal distribution compared to the original data in the RKHS, since the augmentation is a rank $d$ transformation regardless of the size $n$ of the data, which can span an $n$-dimensional subspace in the RKHS (e.g. for universal kernels). Our results show that good representations can be learned with just one non-trivial augmentation that does not resemble the original data. This provides a first theoretical justification for recent empirical observations (Moutakanni et al., 2024).
The Impact of Architecture.
Figure 3 shows that for different architectures, the augmentations are recognizably different. This follows directly from our theory since the matrix $K$ differs depending on the kernel that determines the learnable function class in our setup. It was first noted by Saunshi et al. (2022) that architecture has an implicit bias on the representations, and that this affects downstream performance. Our work also sheds light on the architectural influence in SSL, but from a new angle: For the same target representation, the function class heavily influences the optimal augmentations needed to learn these representations.
The choice of $f^{*}$.
At the heart of our work is the function $f^{*}$ that defines the target representations. While we choose $f^{*}$ to be a pretrained model in our experiments, there is no general consensus on what constitutes a good representation. Our theoretical framework however allows investigating the interplay between augmentation and learned representation from a new angle. For example, suppose a statistical model is assumed for the data-generating process, such as the spiked covariance model that is commonly studied in high-dimensional statistics (Wainwright, 2019):
$x\sim\mathcal{N}_{m}(0,\Sigma)$
$\Sigma=\nu\cdot\theta\theta^{\top}+I_{m}$
where $\nu>0$ and $\theta\in\mathbb{R}^{m}$ is the (unknown) signal direction hidden in the covariance $\Sigma$. Previous works on self-supervised learning would only be able to answer what representations $\hat{\theta}^{\top}x$ are learned from a given augmentation. In contrast, our theoretical results from Section 4 give us analytical expressions for what the “correct” augmentation is, i.e. the one that projects onto the signal direction $\theta$. We leave this for future work to explore.
Acknowledgments
This work was done as part of Shlomo Libo Feigin’s master’s thesis at the Technical University of Munich. It was partially supported by the German Research Foundation (DFG) through the Research Grant GH 257/4-1, and by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research (BMBF). We thank Vladimir Cherkassky, Daniel He, Daniel Bin Schmid, Nimrod de la Vega and Maedeh Zarvandi for the interesting and helpful discussions on this work. We also wish to thank Nil Ayday, Alexandru Craciun and Pascal Esser for the useful comments on the paper.
Impact Statement
This paper furthers our theoretical understanding of the critical role of data augmentations in self-supervised learning. Our work is foundational in nature. While there are potential societal consequences of our work, we do not feel any must be specifically highlighted here. Nonetheless, we point out that the notion of a desired target representation has the potential to reflect the biases of the researcher or ML engineer.
Appendix A Definition of the VICReg Loss
Although in Section 3 we use the variance in the definition of $v$ as part of the VICReg loss, we can use a variaty of other definitions for our results. For example, we can use the standard deviation similar to the original VICReg definition (Bardes et al., 2022). The part of the proof where the definition of VICReg comes to play is Lemma B.2, where we use the fact that:
$v(Z)=0\iff\forall_{i\in[d]}\text{cov}(Z)_{i,i}=1$
Which is of course also true for:
$v(Z)=\frac{1}{d}\sum_{i=1}^{d}\left(1-\sqrt{[\text{cov}\left(Z\right)]_{i,i}}\right)^{2}$
Similar to the definition in Bardes et al. (2022), except we use the $L_{2}$ loss instead of the hinge loss. In fact, any loss function $l(\cdot)$ where $l(y)=0\iff y=0$ would work for our results.
Appendix B Proofs of the Main Results
In addition to the notation in Section 3, we denote $\phi_{i}\coloneqq\phi(x_{i})$ and $\|\cdot\|_{\text{HS}}$ to be the Hilbert-Schmidt norm.
B.1 Proof of Theorem 4.2 (VICReg)
To prove Theorem 4.2 we need a number of auxiliary results. First, we derive a representer-like theorem that guarantees any $\hat{f}$ that minimizes VICReg with least norm to be of the form $\hat{f}=A\Phi^{T}$, provided a suitable augmentation.
Proposition B.1.
Let $\mathcal{L}$ be a joint embedding loss as described in Section 3.1. Let $\mathcal{T}$ be a distribution of augmentations such that $\forall_{T\in\text{Support}(\mathcal{T})}T\Phi\in\{\Phi AK:A\in\mathbb{R}^{n\times n}\}$. Suppose $\hat{f}$ satisfies $\|\hat{f}\|_{HS}=\min\{\|f\|_{HS}:\mathcal{L}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},f)=0\}$ and ($\mathcal{L}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},\hat{f})=0$. Then, $\hat{f}=A\Phi^{T}$ for some $A\in\mathbb{R}^{d\times n}$.
Proof.
Let $\mathcal{T}$ be a distribution of transformations s.t. $\forall_{T\in\text{Support}(\mathcal{T})}T\Phi\in\{\Phi AK:A\in\mathbb{R}^{n\times n}\}$ and let $f\in\mathbb{R}^{d}\otimes\mathcal{H}$ be a solution to ($\mathcal{L}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},f)=0$ Define $S_{\parallel}\coloneqq\{A\Phi^{T}:A\in\mathbb{R}^{d\times n}\}$. We decompose $f=f_{\parallel}+f_{\perp}$ where $f_{\parallel}\in S_{\parallel}$ and $<f_{\perp},\hat{f}>_{\text{HS}}=0$ for every $\hat{f}\in S_{\parallel}$. Since $S_{\perp}$ is finite-dimensional, such decomposition exists via the Gram-Schmidt process.
$\mathcal{L}$ is a function of $\{f(T(\phi_{i}):T\in\text{Support}(\mathcal{T}),i\in[n]\}$, since $\forall_{T\in\text{Support}(\mathcal{T})}T\Phi\in\{\Phi AK:A\in\mathbb{R}^{n\times n}\}$, $\mathcal{L}$ is a function of $\{f(\phi_{i})\}_{i=1}^{n}$.
We first prove that $\forall_{i\in[n]}f_{\perp}(\phi_{i})=0$. Let $A\in\mathbb{R}^{d\times n}$ be an arbitrary matrix.
$0=<f_{\perp},A\Phi^{T}>_{\text{HS}}=Tr(f_{\perp}^{T}A\Phi^{T})=Tr(A\Phi^{T}f_{\perp}^{T})=\sum_{i=1}^{n}<Ae_{i},\Phi^{T}f_{\perp}^{T}e_{i}>$
Where $\Phi^{T}f_{\perp}^{T}$ is a $n\times n$ matrix. Since $A$ is arbitrary it means that $\Phi^{T}f_{\perp}^{T}=0$ and hence $(\Phi^{T}f_{\perp}^{T})^{T}=f_{\perp}\Phi=0$ Therefore: $\forall_{i\in[n]}f_{\perp}(\phi_{i})=0$. We get that for every $i\in[n]$:
$f(\phi_{i})$ $=f_{\parallel}(\phi_{i})$
$\|f\|_{\text{HS}}^{2}$ $=\|f_{\parallel}\|_{\text{HS}}^{2}+\|f_{\perp}\|_{\text{HS}}^{2}$
The min-norm solution of ($\mathcal{L}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},f)=0$ satisfies $f=f_{\parallel}\in S_{\parallel}$.
We continue by deriving two conditions that are equivalent to obtaining zero loss in VICReg.
Lemma B.2.
$L_{VIC}(Z,Z^{\prime})=0$ if and only if $Z=Z^{\prime}$ and $\text{cov}(Z)=n^{-1}ZHH^{T}Z^{T}=I$.
Proof.
$L_{\text{VIC}}(Z,Z^{\prime})=0$ if and only if:
$s(Z,Z^{\prime})$ $=\frac{1}{n}\sum_{i=1}^{n}\left\|z_{i}-z_{i}^{\prime}\right\|_{2}^{2}=0$
$v(Z)$ $=\frac{1}{d}\sum_{i=1}^{d}\left(1-[\text{cov}\left(Z\right)]_{i,i}\right)^{2}=0$
$c(Z)$ $=\frac{1}{d}\sum_{i\neq j}^{d}[\text{cov}(Z)]_{i,j}^{2}=0$
Which is satisfied if and only if $Z=Z^{\prime}$ and $\text{cov}(Z)=I$.
Next, we show a statement that is already very close to Theorem 4.2. The only difference here is that we still assume the target representations to have identity covariance. This simplifies the proof, but will be easy to relax.
Proposition B.3.
Suppose assumption 3.3 is satisfied. Let $\hat{f}\in\mathbb{R}^{d}\otimes\mathcal{H}$ be of the form $\hat{f}=C\Phi^{T}$ s.t. $\text{cov}\left([\hat{f}(\phi_{1}),\ldots,\hat{f}(\phi_{n})]\right)=I_{d}$. Then, any $\tilde{f}$ that minimizes $\mathcal{L}_{VIC}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T}_{\mathcal{H}}^{\text{VIC}}(C),f\right)$ over $\mathbb{R}^{d}\otimes\mathcal{H}$ with least Hilbert-Schmidt-norm, there exists $Q\in O(d)$ s.t. $\tilde{f}=Q\hat{f}$.
Proof.
We first show that $\mathcal{T}_{\mathcal{H}}^{\text{VIC}}(C)$ is well defined by proving that $CKC^{T}$ has full rank. This follows from the fact that if $K$ is a positive definite symmetric matrix, and if $K=LL^{T}$ is the Cholesky decomposition of $K$, $CKC^{T}=(CL)(CL)^{T}$ and $rank(CKC^{T})=rank(CL)=rank(C)=d$.
For simplicity of notation we denote $\mathcal{T}=\mathcal{T}_{\mathcal{H}}^{\text{VIC}}(C)$. Define:
$\mathcal{F}^{\ast}\coloneqq\arg\min_{f\in\mathbb{R}^{d}\otimes\mathcal{H}}\{\|f\|_{\text{HS}}:\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$
We will prove that =$\mathcal{F}^{*}=\{QC\Phi^{T}:Q\in O(d)\}$. Since $\mathcal{L}_{\text{VIC}}$ is non-negative, $\mathcal{F}^{\ast}$ is also the set of minimizers of $\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)$ with least Hilbert-Schmidt-norm.
Since $\mathcal{L}_{\text{VIC}}$ is non-negative, $\mathcal{L}_{\text{VIC}}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},f)=0$ if and only if for every $(T_{1},T^{\prime}_{1},\ldots,T_{n},T^{\prime}_{n})$ in the support of $(\mathcal{T})^{\otimes 2n}$
$L_{\text{VIC}}([f(T_{1}(\phi_{1})),\ldots,f(T_{n}(\phi_{n}))],[f(T^{\prime}_{1}(\phi_{1})),\ldots,f(T^{\prime}_{n}(\phi_{n}))])=0$ (1)
Claim 1.
We prove that $\{QC\Phi^{T}:Q\in O(d)\}\subseteq\{f:\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$.
Let $\tilde{f}=QC\Phi^{T}$ where $Q\in O(d)$. We prove that for every $T,T^{\prime}\in\text{Support}(\mathcal{T})$ and for every $i\in[n]$ it holds that $f(T(\phi_{i}))=f(T^{\prime}(\phi_{i}))$. Since $\text{Support}(\mathcal{T})=\{I_{\mathcal{H}},\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}\}$, it is enough to check that the following holds for every $i\in[n]$:
$QC\Phi^{T}I_{\mathcal{H}}\phi_{i}=QC\Phi^{T}(\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T})\phi_{i}$
Which indeed holds since:
$QC\Phi^{T}\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}\phi_{i}=Q\underbrace{CKC^{T}\left(CKC^{T}\right)^{-1}}_{I}C\Phi^{T}\phi_{i}=QC\Phi^{T}\phi_{i}$
Therefore, (1) is exactly:
$L_{\text{VIC}}(QC\Phi^{T}\Phi,QC\Phi^{T}\Phi)=0$
$L_{\text{VIC}}(QCK,QCK)=0$
Which according to Lemma (B.2) holds if $\text{cov}(QCK)=I$, which is true since we assumed that $\text{cov}(CK)=I$ and therefore indeed $L_{\text{VIC}}(QCK,QCK)=0$, (1) holds and Claim 1 is proven.
Claim 2.
We prove that $\mathcal{F}^{\ast}\subseteq\{QC\Phi^{T}:Q\in O(d)\}$.
Let $\tilde{f}\in\mathcal{F}^{\ast}$. We would like to use Proposition B.1, which requires $\forall_{T\in\text{Support}(\mathcal{T})}T\Phi\in\{\Phi\bar{A}K:\bar{A}\in\mathbb{R}^{n\times n}\}$. For $T=\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}$ the condition is satisfied since by definition $K=\Phi^{T}\Phi$. For $T=I_{\mathcal{H}}$ the condition is satisfied for $\bar{A}=K^{-1}$. Therefore, according to Proposition B.1 $\hat{f}=A\Phi^{T}$.
We now use (1). According to Lemma B.2 for every $i\in[n]$ and $T,T^{\prime}\in\text{Support}(\mathcal{T})$:
$A\Phi^{T}T_{i}(\phi_{i})=A\Phi^{T}T_{i}^{\prime}(\phi_{i})$
Taking $T_{i}=I_{\mathcal{H}}$ and $T_{i}^{\prime}=\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}$ results in the following:
$A\Phi^{T}I_{\mathcal{H}}(\phi_{i})=A\Phi^{T}\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}\phi_{i}$
Which holds for every $i\in[n]$, meaning:
$AK=AKC^{T}\left(CKC^{T}\right)^{-1}CK$ (2)
We use Lemma B.2 again for (1) to obtain that for $T_{1}=\ldots=T_{n}=I_{\mathcal{H}}$
$\text{cov}(A\Phi^{T}T_{1}(\phi_{1}),\ldots,A\Phi^{T}T_{n}(\phi_{n}))=\text{cov}(A\Phi^{T}\Phi)=\text{cov}(AK)=I$
$n^{-1}(AKH)(AKH)^{T}=I$
$n^{-1}AKHH^{T}KA^{T}=I$ (3)
In addition, we assumed that $\text{cov}(CK)=I$, Therefore:
$n^{-1}CKHH^{T}KC^{T}=I$ (4)
We substitute (2) into (3) and get:
$n^{-1}AKC^{T}\left(CKC^{T}\right)^{-1}CKHH^{T}KC^{T}\left(CKC^{T}\right)^{-1}CKA^{T}=I$
We substitute (4) into the above equation and get:
$AKC^{T}\left(CKC^{T}\right)^{-1}\left(CKC^{T}\right)^{-1}CKA^{T}=I$
We denote $Q=AKC^{T}\left(CKC^{T}\right)^{-1}$, the above equation means that $Q\in O(d)$. Substituting $Q$ back to (2) we get:
$AK=QCK$
$A=QC$
Therefore, $\tilde{f}\in\{QC\Phi^{T}:Q\in O(d)\}$ and Claim 2 is proven.
Claim 3.
$\{QC\Phi^{T}:Q\in O(d)\}\subseteq\mathcal{F}^{\ast}$.
We assume by way of contradiction that there exists $\tilde{f}=\tilde{Q}C\Phi^{T}\in\{QC\Phi^{T}:Q\in O(d)\}$ such that $\tilde{f}\notin\mathcal{F}$. We know via Claim 1 that $\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},\tilde{f}\right)=0$, therefore there must be a $\tilde{f}^{\prime}\in\mathcal{F}^{\ast}$ such that $\|\tilde{f}^{\prime}\|_{\text{HS}}<\|\tilde{f}\|_{\text{HS}}$. Applying Claim 2 we know that $\tilde{f}^{\prime}=\tilde{Q}^{\prime}C\Phi^{T}$ for $\tilde{Q}^{\prime}\in O(d)$. We get that $\|\tilde{Q}C\Phi^{T}\|_{\text{HS}}<\|\tilde{Q}^{\prime}C\Phi^{T}\|_{\text{HS}}$ for two matrices $\tilde{Q},\tilde{Q}^{\prime}\in O(d)$. Which is a contradiction since for every $U\in O(d)$:
$\|UC\Phi^{T}\|_{\text{HS}}^{2}=Tr((UC\Phi^{T})^{T}(UC\Phi^{T}))=Tr((C\Phi^{T})^{T}U^{T}U(C\Phi^{T}))=Tr((C\Phi^{T})^{T}(C\Phi^{T})=\|C\Phi^{T}\|_{\text{HS}}^{2}$
Therefore, we have proven Claim 3. Combining Claims 2 and 3 we get $\{QC\Phi^{T}:Q\in O(d)\}=\mathcal{F}^{\ast}$. ∎
With this proposition in hand, Theorem 4.2 is implied directly from the following Proposition.
See 4.2
Proof.
We simply note that if the target representations, collected in a matrix $F^{*}\in\mathbb{R}^{d\times n}$, have a full rank covariance (as guaranteed by Condition 3.2), then there certainly exists an invertible matrix $W\in\mathbb{R}^{d\times d}$ such that the representations $\tilde{F}=WF^{*}$ have identity covariance. From there, Proposition 4.2 applies, choosing $C=\tilde{F}K^{-1}$. Thus, we know that for any least-norm minimizer $\hat{F}$ of VICReg w.r.t the augmnetation $\mathcal{T}^{VIC}_{\mathcal{H}}(C)$, it holds that $WF^{*}=:\tilde{F}=Q\hat{F}$ for some orthogonal $Q$. This implies that $F^{*}=M\hat{F}$ where $M:=W^{-1}Q$ is invertible as the product of two invertible matrices.
B.2 Proof of Theorem 4.4 (Barlow Twins)
We structure the proof by first proving a set of auxiliary results and then proving Theorem 4.4.
Lemma B.4.
Let $T:\mathcal{H}\rightarrow\mathcal{H}$, and $\mathcal{T}$ a distribution of that yields $T$ with probability $0.5$ and $I_{\mathcal{H}}$ with probability $0.5$. $\mathcal{L}_{BT}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{T})=0$ if and only if $(2n)^{-1}A\Phi^{T}(T\Phi\Phi^{T}+\Phi\Phi^{T}T^{T})\Phi A^{T}=I_{d}$
Proof.
Since $\mathcal{L}_{\text{BT}}$ is non-negative, $\mathcal{L}_{\text{BT}}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{T})=0$ if and only if for every $(T_{1},T^{\prime}_{1},\ldots,T_{n},T^{\prime}_{n})$ in the support of $(\mathcal{T}\otimes\mathcal{T}\mid T\neq T^{\prime})^{\otimes n}$:
$L_{\text{BT}}([A\Phi^{T}T_{1}(\phi_{1}),\ldots,A\Phi^{T}T_{n}(\phi_{n})],[A\Phi^{T}T^{\prime}_{1}(\phi_{1}),\ldots,A\Phi^{T}T^{\prime}_{n}(\phi_{n})])=0$
Recall the definition of $L_{\text{BT}}$:
$L_{\text{BT}}(Z,Z^{\prime})$ $=\sum_{i}^{n}(1-\mathcal{C}_{ii})^{2}+\lambda\sum_{i}^{n}(1-\mathcal{C}_{ii})^{2}$
Since $\lambda>0$, we get that $L_{\text{BT}}(Z,Z^{\prime})=0$ if and only if $\mathcal{C}\coloneqq\frac{1}{2n}(ZZ^{\prime T}+Z^{\prime}Z^{T})=I$. Plugging $ZZ^{\prime T}=\sum_{i=1}^{n}(A\Phi^{T}T_{i}(\phi_{1}))(A\Phi^{T}T_{i}^{\prime}(\phi_{1}))^{T}$:
$\frac{1}{2n}\sum_{i=1}^{n}(A\Phi^{T}T_{i}(\phi_{i}))(A\Phi^{T}T_{i}^{\prime}(\phi_{i}))^{T}+(A\Phi^{T}T^{\prime}_{i}(\phi_{i}))(A\Phi^{T}T_{i}(\phi_{i}))^{T}=I$ (5)
Since $\mathcal{T}$ has two augmentations in its support and we draw the augmentations conditioned on $T_{i}\neq T^{\prime}_{i}$, every summand of (5) will be of the following form:
$(A\Phi^{T}T(\phi_{i}))(A\Phi^{T}I_{\mathcal{H}}(\phi_{i}))^{T}+(A\Phi^{T}I_{\mathcal{H}}(\phi_{i}))(A\Phi^{T}T(\phi_{i}))^{T}$
Hence, we write (5) in the following way:
$\frac{1}{2n}\sum_{i=1}^{n}(A\Phi^{T}T(\phi_{i}))(A\Phi^{T}I_{\mathcal{H}}(\phi_{i}))^{T}+(A\Phi^{T}I_{\mathcal{H}}(\phi_{i}))(A\Phi^{T}T(\phi_{i}))^{T}$ $=I$
$\frac{1}{2n}\sum_{i=1}^{n}(A\Phi^{T}T\phi_{i})(A\Phi^{T}\phi_{i})^{T}+(A\Phi^{T}\phi_{i})(A\Phi^{T}T\phi_{i})^{T}$ $=I$
$\frac{1}{2n}\sum_{i=1}^{n}A\Phi^{T}T\phi_{i}\phi_{i}^{T}\Phi A^{T}+A\Phi^{T}\phi_{i}\phi_{i}^{T}T^{T}\Phi A^{T}$ $=I$
$\frac{1}{2n}A\Phi^{T}(T(\sum_{i=1}^{n}\phi_{i}\phi_{i}^{T})+(\sum_{i=1}^{n}\phi_{i}\phi_{i}^{T})T^{T})\Phi A^{T}$ $=I$
$\frac{1}{2n}A\Phi^{T}(T\Phi\Phi^{T}+\Phi\Phi^{T}T^{T})\Phi A^{T}$ $=I$
Lemma B.5.
(See also Simon et al., 2023, Proposition 4.3) Let $\Gamma\in\mathbb{R}^{d\times d}$ be a symmetric positive definite matrix. Let $\Gamma=USU^{T}$ its spectral decomposition. The set of solutions of $W\Gamma W^{T}=I$ is exactly $W=QS^{-\frac{1}{2}}U^{T}$ where $Q\in O(d)$.
Proof.
It is easy to verify that $W=QS^{-\frac{1}{2}}U^{T}$ where $Q\in O(d)$ indeed satisfies $W\Gamma W^{T}=I$. We would like to prove that any $W$ such that $W\Gamma W^{T}=I$ is of that form. Denote $\Gamma^{\frac{1}{2}}=US^{\frac{1}{2}}U^{T}$.
$\Gamma=(\Gamma^{\frac{1}{2}})(U\Gamma^{\frac{1}{2}})^{T}$
$(W\Gamma^{\frac{1}{2}})(W\Gamma^{\frac{1}{2}})^{T}=I$
Hence, $W\Gamma^{\frac{1}{2}}\in O(d)$. Finally:
$W=W\Gamma^{\frac{1}{2}}(\Gamma^{\frac{1}{2}})^{-1}=W\Gamma^{\frac{1}{2}}US^{-\frac{1}{2}}U^{T}$
Where $W\Gamma^{\frac{1}{2}}U\in O(d)$.
Lemma B.6.
Let $C$ be a rank $d$ matrix and $K$ be a positive definite matrix. The set of solutions $A$ to (7) is $\{QC:Q\in O(d)\}$
Proof.
Recall (7):
$AKC^{T}(CKC^{T})^{-2}CKA^{T}=I_{d}$
We denote $\Gamma=(CKC^{T})^{-2}$. First, we show that the least Frobenius norm solutions of $W\Gamma W^{T}$ are $W=QCKC^{T}$.
$CKC^{T}$ is a positive definite and symmetric, and $rank(CKC^{T})=d$. This follows from the fact that if $K$ is a positive definite symmetric matrix, and if $K=LL^{T}$ is the Cholesky decomposition of $K$, $CKC^{T}=(CL)(CL)^{T}$ and $rank(CKC^{T})=rank(CL)=rank(C)=d$.
Let $V\Sigma V^{T}$ be the spectral decomposition of $CKC^{T}$.
$(CKC^{T})^{-2}=V\Sigma^{-2}V^{T}$
Therefore, according to Lemma B.5 the least Frobenius norm solutions of $W\Gamma W^{T}$ are exactly $W=U\Sigma V^{T}$.
$W=U\Sigma V^{T}=UI_{d}\Sigma V^{T}=UV^{T}V\Sigma V^{T}=(UV^{T})CKC^{T}.$
Where $UV^{T}\in O(d)$. Consequently, all solutions to (7) are of the form $A=QC$ where $Q\in O(d)$.
See 4.4
Proof.
For simplicity of notation, we denote $\mathcal{T}=\mathcal{T}_{\mathcal{H}}^{\text{BT}}(C)$. Define:
$\mathcal{F}^{\ast}\coloneqq\arg\min_{f\in\mathbb{R}^{d}\otimes\mathcal{H}}\{\|f\|_{\text{HS}}:\mathcal{L}_{\text{BT}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$
We will prove that =$\mathcal{F}^{*}=\{QC\Phi^{T}:Q\in O(d)\}$.
We begin by rewriting the Lyapunov equation in definition 4.3:
$BK+KB$ $=2n\Psi^{T}\Phi K^{-1}GK^{-1}\Phi^{T}\Psi$
$K$ $=\Psi^{T}\Phi\Phi^{T}\Psi$
$\Psi$ $\coloneqq\Phi K^{-\frac{1}{2}}$
$G$ $\coloneqq KC^{T}(CKC^{T})^{-2}CK$
Notice that $\Phi=\Psi\Psi^{T}\Phi$. Multiplying from the left by $\Psi$ and from the right by $\Psi^{T}$, we get that if $B$ is the solution of the Lyapunov equation, then:
$BK+KB$ $=2n\Psi^{T}\Phi K^{-1}GK^{-1}\Phi^{T}\Psi$
$\Psi BK\Psi^{T}+\Psi KB\Psi^{T}$ $=2n\Psi\Psi^{T}\Phi K^{-1}GK^{-1}\Phi^{T}\Psi\Psi^{T}$
$\Psi B\Psi^{T}\Phi\Phi^{T}\Psi\Psi^{T}+\Psi\Psi^{T}\Phi\Phi^{T}\Psi B\Psi^{T}$ $=2n\Psi\Psi^{T}\Phi K^{-1}GK^{-1}\Phi^{T}\Psi\Psi^{T}$
$\Psi B\Psi^{T}\Phi\Phi^{T}+\Phi\Phi^{T}\Psi B\Psi^{T}$ $=2n\Phi K^{-1}GK^{-1}\Phi^{T}$
$B_{\Phi}\Phi\Phi^{T}+\Phi\Phi^{T}B_{\Phi}$ $=2n\Phi K^{-1}GK^{-1}\Phi^{T}$ (6)
Where $B_{\Phi}\coloneqq\Psi B\Psi^{T}$ is a symmetric operator.
Recall Lemma B.4: See B.4
Meaning that by Definition 4.3:
$\mathcal{L}_{BT}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{T})=0\iff$ $(2n)^{-1}A\Phi^{T}(\Phi K^{-\frac{1}{2}}BK^{-\frac{1}{2}}\Phi^{T}\Phi\Phi^{T}+\Phi\Phi^{T}\Phi K^{-\frac{1}{2}}BK^{-\frac{1}{2}}\Phi^{T})\Phi A^{T}=I_{d}$
$\iff$ $(2n)^{-1}A\Phi^{T}(\Psi B\Psi^{T}\Phi\Phi^{T}+\Phi\Phi^{T}\Psi B\Psi^{T})\Phi A^{T}=I_{d}$
Combining with (6) and the definition of $G$ we get that $A$ is a solution of $\mathcal{L}_{\text{BT}}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{T})=0$ if and only if
$AKC^{T}(CKC^{T})^{-2}CKA^{T}=I_{d}$ (7)
Recall Lemma B.6: See B.6
The following claims directly follow:
Claim 4.
$\{QC\Phi^{T}:Q\in O(d)\}\subseteq\{f:\mathcal{L}_{\text{BT}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$.
And that:
$\arg\min_{f\in\mathbb{R}^{d}\otimes\mathcal{H}}\{\|f\|_{\text{HS}}:\mathcal{L}_{\text{BT}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0,\exists_{A\in\mathbb{R}^{d\times n}}f=A\Phi^{T}\}\subseteq\{QC\Phi^{T}:Q\in O(d)\}$
Proposition B.1 implies that:
=$\mathcal{F}^{\ast}=\{\|f\|_{\text{HS}}:\mathcal{L}_{\text{BT}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0,\exists_{A\in\mathbb{R}^{d\times n}}f=A\Phi^{T}\}$
Therefore:
Claim 5.
$\mathcal{F}^{\ast}\subseteq\{QC\Phi^{T}:Q\in O(d)\}$.
Claim 6.
$\{QC\Phi^{T}:Q\in O(d)\}\subseteq\mathcal{F}^{\ast}$.
We assume by way of contradiction that there exists $\tilde{f}=\tilde{Q}C\Phi^{T}\in\{QC\Phi^{T}:Q\in O(d)\}$ such that $\tilde{f}\notin\mathcal{F}$. We know via Claim 4 that $\mathcal{L}_{\text{BT}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},\tilde{f}\right)=0$, therefore there must be a $\tilde{f}^{\prime}\in\mathcal{F}^{\ast}$ such that $\|\tilde{f}^{\prime}\|_{\text{HS}}<\|\tilde{f}\|_{\text{HS}}$. Applying Claim 5 we know that $\tilde{f}^{\prime}=\tilde{Q}^{\prime}C\Phi^{T}$ for $\tilde{Q}^{\prime}\in O(d)$. We get that $\|\tilde{Q}C\Phi^{T}\|_{\text{HS}}<\|\tilde{Q}^{\prime}C\Phi^{T}\|_{\text{HS}}$ for two matrices $\tilde{Q},\tilde{Q}^{\prime}\in O(d)$. Which is a contradiction since for every $U\in O(d)$:
$\|UC\Phi^{T}\|_{\text{HS}}^{2}=Tr((UC\Phi^{T})^{T}(UC\Phi^{T}))=Tr((C\Phi^{T})^{T}U^{T}U(C\Phi^{T}))=Tr((C\Phi^{T})^{T}(C\Phi^{T})=\|C\Phi^{T}\|_{\text{HS}}^{2}$
Therefore, we have proven Claim 6. Combining Claims 5 and 6 we get $\{QC\Phi^{T}:Q\in O(d)\}=\mathcal{F}^{\ast}$. ∎
B.3 Proof of Proposition 4.5
See 4.5
Proof.
The result is a corollary of Theorems 4.2 and 4.4, as they imply that for a least norm minimizer of ($\mathcal{L}\Bigl{(}\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(FK^{-1}),\;f\Bigr{)}$ there exist an invertible matrix $M\in\mathbb{R}^{d\times d}$ and a vector $b\in\mathbb{R}^{d}$ such that
$f=MFK^{-1}\Phi^{T}+b$
Implying:
$f\Phi$ $=MFK^{-1}\Phi^{T}\Phi+b$
$f\Phi$ $=MFK^{-1}K+b$
$f\Phi$ $=MF+b$
$\forall_{i\in[n]}f(\phi(x_{i}))$ $=Mf^{*}(x_{i})+b$
Appendix C Details on the Experiments
For all of the experiments we used a pretrained ResNet given by ResNet50_Weights.IMAGENET1K_V2 in PyTorch (Paszke et al., 2019). we reduced the dimension of the target representations to 64 using PCA.
For calculating the augmentations in Figure 3 we used 10,000 images of MNIST training data as $\{x_{i}\}_{i=1}^{n}$ and we calculated the augmented images of validation data which was not part of $\{x_{i}\}_{i=1}^{n}$ in Algorithm 1, meaning these are augmentations produced for new images outside of the “training set”. We used $\sigma=3$ for the RBF kernel and $\lambda_{\text{ridge}}=1,\mu_{P}=1$ ($\mu_{P}$ being a parameter for solving the preimage problem) for these experiments.
For Figure 4 we used the digits 0 and 1 of a training set consisting of 50,000 MNIST images. We used the RBF kernel with $\sigma=1$ and $\lambda=5,\mu=5,\nu=1$ for the VICReg losses. We optimized the loss as one batch consisting of the training images and their augmented views. We parameterized the learnable function as $f=C\Phi^{T}$ where $C\in\mathbb{R}^{d\times n}$ are the learnable parameters, the calculations were done using the kernel trick. We used the Adam optimizer (Kingma & Ba, 2017) with a learning rate of 0.001.