### v4

# A Theoretical Characterization of Optimal Data Augmentations in Self-Supervised Learning

Abstract
Data augmentations play an important role in the recent success of self-supervised learning (SSL). While augmentations are commonly understood to encode invariances between different views into the learned representations, this interpretation overlooks the impact of the pretraining architecture and suggests that SSL would require diverse augmentations which resemble the data to work well. However, these assumptions do not align with empirical evidence, encouraging further theoretical understanding to guide the principled design of augmentations in new domains. To this end, we use kernel theory to derive analytical expressions for data augmentations that achieve desired target representations after pretraining. We consider non-contrastive and contrastive losses, namely VICReg, Barlow Twins and the Spectral Contrastive Loss, and provide an algorithm to construct such augmentations. Our analysis shows that augmentations need not be similar to the data to learn useful representations, nor be diverse, and that the architecture has a significant impact on the optimal augmentations.
1 Introduction
Self-supervised learning (SSL) has gained prominence in recent years, serving as one of the backbones of the foundation models driving current progress in artificial intelligence. Instead of using labels, SSL employs a surrogate objective to learn representations, which are then used for downstream tasks. Joint embedding methods (Chen et al., 2020b; He et al., 2020; Grill et al., 2020; Zbontar et al., 2021) in particular have seen a rise in popularity, achieving competitive performance with supervised representation learning for downstream classification, image segmentation and object detection (Chen and He, 2020; Bardes et al., 2022). Broadly speaking, these models encourage different views of the same underlying point to align closely in the embedding space, while preventing the representations from collapsing (Jing et al., 2022). In the vision domain, the views are usually given by augmentations such as random crop, Gaussian blur, and color distortion (Chen et al., 2020b).
The choice of augmentations is critical to the representations learned by the SSL objective. Different downstream tasks may require different augmentation; for example, Purushwalkam and Gupta (2020) show that cropping encourages invariance to occlusions but negatively affects downstream tasks that require category and viewpoint invariance. In addition, Xiao et al. (2020); Lee et al. (2021); Zhang and Ma (2022); Bendidi et al. (2023) demonstrate that certain augmentations benefit or penalize downstream classification on different classes, and Ericsson et al. (2022) empirically show that pose-related tasks and classification-related tasks benefit from opposite augmentations. Previous theoretical studies largely do not address the subtleties of augmentation choice. Instead, by assuming a certain relationship between the data and the augmentations, they provide guarantees specifically on downstream classification tasks (Arora et al., 2019; HaoChen et al., 2022; Saunshi et al., 2022).
In practice, however, the choice of suitable augmentations depends not only on the downstream task, but also intricately on the domain (Bendidi et al., 2023; Balestriero et al., 2023). Medicine still struggles to benefit from the success of SSL, arguably since assumptions that hold for natural images do not hold for medical images (Huang et al., 2023). As an example, consider brain scans. Unlike natural images, which typically include a central object, cropping parts of an image with a tumor can drastically change the interpretation of the scan. In practice, augmentations are therefore carefully crafted for different applications, a manual and empirically driven process. Theoretical insights are scarce, and as we discuss next, the few that exist do not explain empirical phenomena well. For example, if augmentations primarily serve the role of generating different views, one would expect that data augmentations need to (a) be similar to the original data, and (b) capture a diverse set of views. Neither holds in practice. Firstly, strong data augmentations, such as random crop and cutout, have drastically different marginal distributions than the original data (Gontijo-Lopes et al., 2020). At the same time, these are exactly the augmentations that empirically provide the most benefit for downstream performance (Chen et al., 2020b). In contrast, natural-seeming augmentations, such as adding Gaussian noise, only show limited benefit. Secondly, Bardes et al. (2022); Cabannes et al. (2023) find that it is always better to use more data instead of more augmentations. In fact, Chen et al. (2020b) use only three types of augmentations. Finally, Moutakanni et al. (2024) stretch both of the above assumptions to the limit, showing that with just one augmentation, cropping, one can achieve state-of-the-art representations provided sufficient data.
(a) Previous works
(b) Our study
(c) Our approach
Figure 1: (a) Previous theoretical works assume certain augmentation characteristics and examine the learned representations.  (b) Our study asks the reverse question: given target representations (e.g. from a pretrained ResNet), what augmentations achieve them?  (c) An illustration of our pipeline: we find a transformation $T_{\mathcal{H}}$ in the RKHS that yields the target representations $f^{*}$ (Theorems 4.24.34.5), then translate $T_{\mathcal{H}}$ back to input space by solving a pre-image problem $\tilde{\phi}^{-1}$ (Algorithm 1).
In self-supervised learning, a unique interplay exists between data, augmentations and representations (Cabannes et al., 2023). Prior works predominantly focus on one side of this interaction, asking how data and augmentations influence the learned representations. In this paper, we study SSL from a different angle. We ask: For given data and desired target representations, what augmentations result in these representations? This conceptually deviates from prior studies, as illustrated in Figure 1.
Contributions.
To tackle this question, we place ourselves within existing theoretical frameworks for SSL. We formally prove that suitable data augmentations together with a sufficiently expressive hypothesis class can learn any desired representation by joint embedding methods such as VICReg (Bardes et al., 2022), Barlow Twins (Zbontar et al., 2021) and the Spectral Contrastive Loss (HaoChen et al., 2022) — the latter being a theoretical proxy to SimCLR (Chen et al., 2020b). We derive analytical expressions for the augmentations in Section 4, and propose an algorithm to compute these augmentations in Section 5. Finally, we give insights about augmentation choice in Section 7, interpreting the aforementioned empirical phenomena through the lens of our analysis. To summarize, our main contributions are the following.
  • We prove that for VICReg, Barlow Twins and the Spectral Contrastive Loss, it is possible to guarantee the recovery of any representation for the input data, given suitable augmentations and a sufficiently expressive function class.
  • For VICReg and the Spectral Contrastive Loss, we derive a closed-form solution for the augmentations; for Barlow Twins, the augmentations are expressed through a solution of a continuous-time Lyapunov equation. To the best of our knowledge, this is the first method to construct augmentations explicitly for any given target representation.
  • Our theoretical results provide new insights into the role of augmentations in SSL: (a) Augmentations need not be similar to the original data. (b) Even very few augmentations can provide good representations, as empirically observed by Moutakanni et al. 2024. (c) Augmentations corresponding to the same representations can be recognizably different depending on the architecture. (d) Augmentations can act as projections in the feature space as opposed to different views of the data.
2 Related Work
Theory of Self-Supervised Learning.
Previous theoretical works on self-supervised learning primarily focus on mathematically describing the learned representations, while implicitly assuming that useful augmentations are already given. This high-level idea has been formalized in numerous ways. Positive pairs can be assumed to be independent samples from the same class (Arora et al., 2019), or a data augmentation graph is defined in which positive pairs are connected (HaoChen et al., 2022). Furthermore, a suitable integral operator can encode similarity between positive pairs (Cabannes et al., 2023), or the target representations are assumed to be contained in a Hilbert space defined by the augmentation (Zhai et al., 2024). In this work, we turn this question around: Given target representations, what are the augmentations one needs to learn these representations?
Kernel Methods and SSL.
Studying supervised deep learning through kernel methods has proven to be fruitful, most notably by virtue of the neural tangent kernel (NTK) (Jacot et al., 2018). This has sparked several works that assume kernel models for the representation function in SSL (Kiani et al., 2022; Balestriero and LeCun, 2022; Simon et al., 2023; Cabannes et al., 2023; Esser et al., 2024). Indeed, the validity of the NTK approximation in SSL has recently been proven for Barlow Twins (Fleissner et al., 2024), justifying the use of kernel theory to understand SSL with neural networks. Additionally, it is known that contrastive learning can in fact be viewed as kernel learning (Johnson et al., 2022), which has been used to derive new generalization error bounds for SSL (Zhai et al., 2024). In this paper, we therefore also consider a kernel setting. For VICReg our framework coincides with that of Balestriero and LeCun (2022), for Barlow Twins our framework matches Simon et al. (2023), and for the Spectral Contrastive Loss our framework is similar to that of (Esser et al., 2024).
The Role of Data Augmentations.
The role of data augmentations has been studied both in the supervised setting as well as in the self-supervised setting. In the supervised setting, Gontijo-Lopes et al. (2020); Kim et al. (2022) suggest that augmentations should be similar to the original data distribution, and diverse. Chen et al. (2020a) view augmentations as group actions that keep the data distribution approximately similar. Geiping et al. (2023) show that augmentations can be helpful even if they encode invariances not present in the original data. In SSL, Wang and Isola (2020) explain the role of data augmentations through the alignment of the positive samples and the uniformity of induced distribution. Wang et al. (2022) present the role of augmentations as connecting the data through overlaps, which then cause class-separated representations in downstream classification. Kügelgen et al. (2021) view augmentations as transformations that preserve the semantic content of an image while changing the style component. Taking an information theoretic perspective, Tian et al. (2020) argue that augmentations should reduce the mutual information between views as much as possible while keeping the shared task-relevant information intact. It is challenging (and in some cases even impossible) to reconcile these interpretations with the empirical observation that augmentations produce useful representations even when they are not diverse or similar to the original distribution (Moutakanni et al., 2024).
3 Preliminaries
3.1 Problem Statement and Approach
Joint Embedding Loss Functions.
Throughout this paper, we assume we are given $n$ data points $\{x_{i}\}_{i=1}^{n}\subseteq\mathcal{X}\subseteq\mathbb{R}^{m}$. For each of these points $x_{i}$, we create two views by sampling from a random augmentation map $T:\mathcal{X}\rightarrow\mathcal{X}$ with distribution $\mathcal{T}$. This yields $T_{i}(x_{i})$ and $T^{\prime}_{i}(x_{i})$. Both views are passed to a function $f:\mathcal{X}\rightarrow\mathbb{R}^{d}$ that maps to a lower-dimensional space with $d<\min(n,m)$. In practice, $f$ is typically a neural network, while theoretical works often consider $f$ to be a kernel function (Kiani et al., 2022; Balestriero and LeCun, 2022; Simon et al., 2023; Cabannes et al., 2023; Esser et al., 2024). We denote $\mathcal{F}$ for the function class from which $f$ is chosen. Denoting $z_{i}=f(T(x_{i}))$ and $z_{i}^{\prime}=f(T^{\prime}(x_{i}))$, a joint embedding loss $L(Z,Z^{\prime})$ is computed on the matrices $Z=[z_{1},\ldots,z_{n}]$ and $Z^{\prime}=[z_{1}^{\prime},\ldots,z_{n}^{\prime}]\in\mathbb{R}^{d\times n}$. The loss minimized over $\mathcal{F}$ is of the form ($\mathcal{L}\Bigl{(}\{x_{i}\}_{i=1}^{n},\mathcal{T},f\Bigl{)}=\mathbb{E}_{Z,Z^{\prime}}\Bigl{[}L(Z,Z^{\prime})\Bigl{]}$, where the expectation is with respect to the randomness in the augmentations. Depending on the method, $T_{i}$ and $T^{\prime}_{i}$ can either be independent samples from the underlying distribution $\mathcal{T}$, or they can be conditioned on the event $T^{\prime}_{i}\neq T_{i}$ (in that case, augmentations always yield two distinct views of the same point $x_{i}$). .
Kernels, RKHS and Function class.
We assume that we are given a kernel $\kappa:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. The canonical feature map of this kernel $\kappa$ is denoted by $\phi(x):\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is the reproducing kernel Hilbert space (RKHS) associated with $\kappa$. From there, we use the tensor product notation $\mathbb{R}^{d}\otimes\mathcal{H}$ to denote the set of linear maps from $\mathcal{H}$ to $\mathbb{R}^{d}$. An element $\Theta\in\mathbb{R}^{d}\otimes\mathcal{H}$ can be thought of as a matrix with $d$ rows that are elements of $\mathcal{H}$. The adjoint operator, denoted as $\Theta^{\top}\in\mathcal{H}\otimes\mathbb{R}^{d}$, can be thought of as the transposed matrix. We equip the function space with the Hilbert-Schmidt norm on $\mathbb{R}^{d}\otimes\mathcal{H}$. The space $\mathcal{F}$ from which our model $f$ is chosen is given by the set of all functions $x\mapsto\Theta^{\top}\phi(x)$ where $\Theta\in\mathbb{R}^{d}\otimes\mathcal{H}$, again equipped with the Hilbert-Schmidt norm on $\Theta$. In this paper, we restrict ourselves to the least norm solutions of ($\mathcal{L}\left(\{\phi(x_{i})\}_{i=1}^{n},\mathcal{T},f\right)$ on $\mathcal{F}$, as is commonly done in kernel methods. Formally:
Definition 3.1 (Least Norm Minimizers).
A function $\hat{f}$ is said to minimize a loss $\mathcal{L}$ with least norm over $\mathcal{F}$ if $\mathcal{L}(\hat{f})=\inf\limits_{f\in\mathcal{F}}\mathcal{L}(f)$ and $\|\hat{f}\|_{\mathcal{F}}\leq\|f\|_{\mathcal{F}}$ for all $f$ with $\mathcal{L}(f)=\inf\limits_{g\in\mathcal{F}}\mathcal{L}(g)$.
Problem Statement.
Given data $\{x_{i}\}_{i=1}^{n}\subset\mathcal{X}$ and a desired target function $f^{\ast}:\mathcal{X}\rightarrow\mathbb{R}^{d}$, our goal is to find a distribution $\mathcal{T}$ of random transformations such that $f^{*}$ is equivalent to $\min_{f\in\mathcal{F}}\mathcal{L}(\{x_{i}\}_{i=1}^{n},\mathcal{T},f)$. We consider two functions $f$ and $g$ to be equivalent if they are identical up to an invertible affine transformation. This notion of equivalence is justified by HaoChen et al. (2022, Lemma 3.1), who prove that invertible affine transformations do not influence the downstream performance when using the standard linear evaluation protocol (Chen et al., 2020b).
Motivated by the empirical study of Moutakanni et al. (2024), we consider distributions $\mathcal{T}$ that can lead to only two possible augmentations — the identity and some other transformation $T$. We construct $T$ by first finding a suitable augmentation in the Hilbert space $\mathcal{H}$, and only later translating it back to the input space $\mathcal{X}$. Given a target representation $f^{*}$, we first identify a binary distribution $\mathcal{T}_{\mathcal{H}}$ that either samples from the identity operator in $\mathcal{H}$ or reduces to a transformation $T_{\mathcal{H}}$. The distribution $\mathcal{T}_{\mathcal{H}}$ is carefully chosen in a way that ensures the least-norm minimizer of the loss is equivalent to the desired target $f^{*}$. Having found the correct augmentation in the Hilbert space, we translate it back to the input space by defining $T\coloneqq\tilde{\phi}^{-1}\circ T_{\mathcal{H}}\circ\phi$ where $\tilde{\phi}^{-1}$ is a solution to the pre-image problem for kernel machines. Our approach is illustrated in Figure 1(c).
Continuous-Time Lyapunov Equations.
Our theoretical results for the Barlow Twins loss partly build on solutions to continuous-time Lyapunov equations. These are matrix equations of the form $AX+XA^{\top}=C$, where $X,A,C$ are matrices of appropriate size. For our purposes, the main result we need is that if $C$ is symmetric and $A$ is a positive-definite matrix, then there is a unique symmetric solution $X$ that solves the Lyapunov equation (Ortega, 1987, Theorem 6.4.2). The solution $X$ can be represented analytically in terms of $C$ and $A$.
Additional Notation.
[…]
Assumptions.
In this paper, we assume two conditions to be satisfied. Firstly, the target representations should have full rank. This condition is reasonably mild: if $f^{*}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$ has linearly dependent dimensions in its range, then there exists an equivalent $f^{\prime}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d^{\prime}}$ with independent dimensions such that $f^{*}$ and $f^{\prime}$ will perform the same under the standard linear evaluation protocol.
Condition 3.2.
Given a target function $f^{*}$ and data $\{x_{i}\}_{i=1}^{n}$, the covariance matrix $\text{cov}\left([f^{\ast}(x_{1}),\ldots,f^{\ast}(x_{n})]\right)$ has full rank $d$.
Secondly, the Gram matrix of the kernel should be invertible. For universal kernels (e.g. the Gaussian RBF kernel), Condition 3.3 holds for any set of distinct samples $\{x_{i}\}_{i=1}^{n}$.
Condition 3.3.
Given a kernel function $\kappa(\cdot,\cdot)$ and data $\{x_{i}\}_{i=1}^{n}$, $K=[\kappa(x_{i},x_{j})]_{i,j}$ has full rank.
3.2 Non-Contrastive and Contrastive Loss Functions
We present the loss functions we analyze. All claims made are proven in the appendix.
VICReg.
The VICReg loss (Bardes et al., 2022) is defined via three components:
$\mathcal{L}_{\text{VIC}}\Bigl{(}\{x_{i}\}_{i=1}^{n},\mathcal{T},f\Bigl{)}=\mathbb{E}_{Z,Z^{\prime}}\Bigl{[}\lambda s(Z,Z^{\prime})+\mu[v(Z)+v(Z^{\prime})]+\nu[c(Z)+c(Z^{\prime})]\Bigl{]}\text{, where}$
$s(Z,Z^{\prime})=\frac{1}{n}\sum_{i=1}^{n}\|z_{i}-z_{i}^{\prime}\|_{2}^{2},\quad v(Z)=\frac{1}{d}\sum_{i=1}^{d}\left(1-[\text{cov}\left(Z\right)]_{i,i}\right)^{2},\quad c(Z)=\frac{1}{d}\sum_{i\neq j}^{d}[\text{cov}(Z)]_{i,j}^{2}$
and $\lambda,\mu,\nu>0$ are hyper-parameters. $z_{i}=f\Bigl{(}T_{i}(x_{i})\Bigl{)},z_{i}^{\prime}=f\Bigl{(}T^{\prime}_{i}(x_{i})\Bigl{)}$, where $\left(T_{i},T^{\prime}_{i}\right){\sim}\mathcal{T}\otimes\mathcal{T}$.
This definition is identical to the original VICReg loss (Zbontar et al., 2021), except we use the $L_{2}$ loss instead of the hinge loss in the definition of $v$. To be consistent with previous theoretical works (Balestriero and LeCun, 2022; Cabannes et al., 2023), we define $v$ using the variance instead of the standard deviation, however our results still hold when using the standard deviation (Appendix A).
Barlow Twins.
The Barlow Twins Loss (Zbontar et al., 2021) is defined in the following way:
$\mathcal{L}_{\text{BT}}\Bigl{(}\{x_{i}\}_{i=1}^{n},\mathcal{T},f\Bigl{)}=\mathbb{E}_{Z,Z^{\prime}}\Bigl{[}\sum_{i}^{n}(1-\mathcal{C}_{ii})^{2}+\lambda\sum_{i\neq j}(1-\mathcal{C}_{ij})^{2}\Bigl{]};\quad\mathcal{C}=\frac{1}{2n}(ZZ^{\prime\top}+Z^{\prime}Z^{\top})$
Where $\lambda>0$ is a hyperparameter. $z_{i}=f\Bigl{(}T_{i}(x_{i})\Bigl{)},z_{i}^{\prime}=f\Bigl{(}T^{\prime}_{i}(x_{i})\Bigl{)}$, where the random augmentations $\left(T_{i},T_{i}^{\prime}\right){\sim}\mathcal{T}\otimes\mathcal{T}$ are conditioned on the event $T_{i}\neq T^{\prime}_{i}$. Similar to Simon et al. (2023), we consider a symmetrized version of the unnormalized cross-correlation for $\mathcal{C}$.
Spectral Contrastive Loss.
The Spectral Contrastive Loss (SCL) (HaoChen et al., 2022) is a theoretical proxy to the SimCLR loss. We define the following sample loss:
$\mathcal{L}_{\text{SCL}}\Bigl{(}\{x_{i}\}_{i=1}^{n},\mathcal{T},f\Bigl{)}$ $=\mathbb{E}_{Z,Z^{\prime}}\Bigl{[}-\frac{2}{n}{}\sum_{i=1}^{n}z_{i}^{\top}z^{\prime}_{i}+\frac{1}{n^{2}}\sum_{i\neq j}(z_{i}^{\top}z^{\prime}_{j})^{2}+\frac{1}{2n^{2}}\sum_{i=1}^{n}(\|z_{i}\|^{4}+\|z_{i}^{\prime}\|^{4})\Bigl{]}$
$z_{i}=f\Bigl{(}T_{i}(x_{i})\Bigl{)},z_{i}^{\prime}=f\Bigl{(}T^{\prime}_{i}(x_{i})\Bigl{)}$, where $\left(T_{i},T^{\prime}_{i}\right){\sim}\mathcal{T}\otimes\mathcal{T}$.
Remark.
Without regularization, the term $-\frac{2}{n}{}\sum_{i=1}^{n}z_{i}^{\top}z^{\prime}_{i}+\frac{1}{n^{2}}\sum_{i\neq j}(z_{i}^{\top}z^{\prime}_{j})^{2}$ alone can diverge to $\infty$- ∞ for small batches, as we show in the appendix. Previous definitions of the sample loss for SCL deal with that fact in various ways. For example, HaoChen et al. (2022) enforce the norm of representations to be fixed in their experiments, while Esser et al. (2024) regularize using the norm $\|\cdot\|_{\mathcal{H}}$. The regularization term $\frac{1}{2n^{2}}\sum_{i=1}^{n}(\|z_{i}\|^{4}+\|z_{i}^{\prime}\|^{4})$ is theoretically motivated by a guaranteed tight bound $\mathcal{L}_{\text{SCL}}\geq-d$ for any $n$, which we prove in the appendix. As $n\rightarrow\infty$, the regularization decreases and $\mathcal{L}_{\text{SCL}}$ uniformly converges to the population SCL as defined in HaoChen et al. (2022).
4 Main Results
In this section, we develop our main results for VICReg (Theorem 4.2), the Spectral Contrastive Loss (Theorem 4.3) and Barlow Twins (Theorem 4.5), proving that any desired target representations can be found by using suitable augmentations, up to equivalence. We defer the proofs to the appendix. We begin by defining the optimal augmentations for VICReg and SCL, and then verify their optimality.
Definition 4.1 (Optimal VICReg and SCL Augmentations).
Consider a target representation of the form $f^{*}=C\Phi^{\top}$, where $C\in\mathbb{R}^{d\times n}$ has full rank. Assuming Condition 3.3, we define $\mathcal{T}_{\mathcal{H}}(C)$ as a distribution of transformations yielding $I_{\mathcal{H}}$ and $\Phi\,C^{\top}\bigl{(}C\,K\,C^{\top}\bigr{)}^{-1}C\,\Phi^{\top}$ with probability $\frac{1}{2}$ each.
Theorem 4.2 (Optimality of Augmentations for VICReg).
[…]
Interestingly, the exact same augmentations can be used for SCL:
Theorem 4.3 (Optimality of Augmentations for SCL).
[…]
Notably, the transformation $T=\Phi\,C^{\top}\bigl{(}C\,K\,C^{\top}\bigr{)}^{-1}C\,\Phi^{\top}$ has rank $d$ and $T^{2}=T$, whereas the data spans a subspace of dimension $n>d$ in $\mathcal{H}$ as a consequence of Condition 3.3. Thus, $T$ can be interpreted as a projection to a low dimensional subspace in the feature space $\mathcal{H}$. In particular, the augmentations induced by $\mathcal{T}_{\mathcal{H}}$ will have a different marginal distribution than the data, no matter what $f^{*}$ is. This is contrary to previous interpretations (discussed in Section 2), which argue that data augmentations should produce views similar to the data to learn a “good” $f^{*}$.
Our results for Barlow Twins take a similar form. We again state the optimal augmentations first.
Definition 4.4 (Optimal Barlow Twins Augmentations).
Given a rank $d$ matrix $C\in\mathbb{R}^{d\times n}$ and assuming Condition 3.3, we define $\mathcal{T}_{\mathcal{H}}^{\text{BT}}(C)$ as a distribution of transformations yielding $I_{\mathcal{H}}$ and $\Phi K^{-\frac{1}{2}}BK^{-\frac{1}{2}}\Phi^{\top}$ with probability $\frac{1}{2}$ each, where $B$ is the unique solution to the continuous-time Lyapunov equation $KB+BK^{\top}=2n\cdot K^{\frac{1}{2}}C^{\top}\Bigl{(}CKC^{\top}\Bigr{)}^{-2}CK^{\frac{1}{2}}$.
Theorem 4.5 (Optimality of Augmentations for Barlow Twins).
[…]
Remark.
Assuming the functional form $f^{*}=C\Phi^{\top}$ for the target representation is not a restrictive condition. By virtue of the representer theorem (Schölkopf et al., 2001), any least norm minimizer of a loss function that only depends on the training data is certainly contained in the span of $\{\phi(x_{i})\}_{i=1}^{n}$. For a general function $f^{*}:~{}\mathcal{X}\rightarrow\mathbb{R}^{d}$, for example a pretrained ResNet, Corollary 4.6 below allows exactly reconstructing representations $\{f^{*}(x_{i})\}_{i=1}^{n}$ on any set of training data.
Corollary 4.6 (Reconstruction of General Representations).
[…]
5 Algorithm
Based on our theoretical results, we provide Algorithm 1 to compute the augmentations in the input space given an arbitrary target $f^{*}$ (e.g. a pretrained ResNet or a ViT). Below, we discuss the preimage problem, matching $f^{*}$ on new data, and how the algorithm can be scaled to large datasets.
Algorithm 1 Augmentation Learning for SSL
[…]
2:Augmented data $\{\hat{x}^{\prime}_{i}\}_{i=1}^{k}$
3:$K\leftarrow$ Kernel matrix on $\{x_{i}\}_{i=1}^{n}$;  $K_{X,\hat{X}}\leftarrow$ Cross-kernel matrix between $\{x_{i}\}_{i=1}^{n}$ and $\{\hat{x}_{i}\}_{i=1}^{k}$
4:$C\leftarrow[f^{*}(x_{1}),\ldots,f^{*}(x_{n})](K+\lambda_{\text{ridge}}I)^{-1}$$\triangleright$ Solve KRR to obtain representer coefficients.
5:if method is VICReg or SCL then
6:     $C_{aug}\leftarrow C^{T}\left(CKC^{T}\right)^{-1}CK_{X,\hat{X}}$triangleright Compute augmentation coefficients (Thms. 4.2, 4.3).
7:end if
8:if method is Barlow Twins then
9:     $B\leftarrow$ Solution of the Lyapunov equation $KB+BK^{\top}=2nK^{\frac{1}{2}}C^{\top}(CKC^{\top})^{-2}CK^{\frac{1}{2}}$
10:     $C_{aug}\leftarrow K^{-\frac{1}{2}}BK^{-\frac{1}{2}}K_{X,\hat{X}}$triangleright Compute augmentation coefficients (Thm. 4.5).
11:end if
12:return solution of the pre-image problem for $\phi^{\prime}=\Phi C_{aug}$.
The Pre-Image Problem in Kernel Machines.
In the previous section, we derived $T_{\mathcal{H}}$ as maps in the RKHS $\mathcal{H}$, taking the form of $\phi(x_{i})\mapsto\Phi M\Phi^{\top}\phi(x_{i})$ for a matrix $M\in\mathbb{R}^{n\times n}$. In particular, this construction yields augmented versions of original points $\phi(x)$ that are of the form $\Phi\theta$ for some $\theta\in\mathbb{R}^{n}$, and as such lie in the Hilbert space. In this section, we translate these augmentations back to the input space by identifying $x^{\prime}\in\mathcal{X}$ such that $\phi(x^{\prime})\approx\Phi\theta$. In general, such $x^{\prime}$ need not exist (Mika et al., 1998) — finding an approximation is a task known as the pre-image problem for kernel machines (Kwok and Tsang, 2004). For our purposes, we use the closed-form approximation proposed by Honeine and Richard (2011), which we detail in the appendix. We emphasize that while our augmentations $T_{\mathcal{H}}$ are formally defined in the Hilbert space, the kernel trick allows all computations to happen in the input space. Moreover, for new data points $x^{\prime}$, the functional form of the augmentations allows evaluating the transformation of $\phi(x^{\prime})$.
Generalization on New Data.
As mentioned in the previous section, our method allows exactly matching any desired target representation $f^{*}$ on given training data $\{x_{i}\}_{i=1}^{n}$ up to equivalence (i.e. up to an invertible affine transformation). It is natural to question the statistical soundness of interpolating on target representations $f^{*}(x_{i})$ obtained e.g. from a ResNet. Therefore, we also incorporate the possibility of using a ridge parameter $\lambda>0$ to avoid overfitting. We solve $\arg\min_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\left\|f^{*}(x_{i})-f(x_{i})\right\|^{2}+\lambda\|f\|_{\mathcal{F}}^{2}$ where $\mathcal{F}$ is, as before, the space of functions of the form $x\mapsto\Theta^{\top}\phi(x),\Theta\in\mathbb{R}^{d}\otimes\mathcal{H}$ equipped with the Hilbert-Schmidt norm on $\Theta$. The optimal solution is $f=F(K+\lambda_{\text{ridge}}I)^{-1}\Phi^{\top}$, where we denote $F=\Bigl{[}f^{*}(x_{1}),\ldots,f^{*}(x_{n})\Bigr{]}$. Crucially, this is still a function of the form $f=C\Phi^{\top}$, and so we can use the augmentations $\mathcal{T}_{\mathcal{H}}\Bigl{(}F(K+\lambda_{\text{ridge}}I)^{-1}\Bigr{)}$ defined in Definition 4.1 and 4.4 respectively to obtain augmentations that achieve the desired target representations — up to some small error that is introduced by $\lambda_{\text{ridge}}>0$.
Scalability.
Much work has been done to make kernel methods scalable. Of particular importance are random feature methods (Rahimi and Recht, 2007) and the Nyström method (Williams and Seeger, 2000). Kernel ridge regression has been adapted to utilize GPU hardware (Meanti et al., 2020, 2022). These methods propose approximators of the form $f=C[\phi(x_{1}),\ldots,\phi(x_{n^{\prime}})]^{\top}$ where where $C$ is calculated as an efficient solution to KRR for $n$ points, $n^{\prime}<n$. These approximate solutions are of the functional form necessitated by Theorems 4.2, 4.3,4.5 and therefore can be used in Algorithm 1.
6 Experiments
Visualizing the Augmentations.
We apply Algorithm 1 to MNIST (LeCun et al., 2010) using representations obtained from a ResNet50 (He et al., 2016) architecture pretrained on ImageNet (Deng et al., 2009). We show the results for VICReg and SCL using different kernel functions in Figure 2(a). It can be observed that the augmented images are dissimilar from the original data, and that different kernel architectures result in semantically different augmentations. This is a direct consequence of the analytical form of the augmentations, which depend on the kernel matrix $K$.
Evaluating the Reconstruction of $f^{*}$.
To verify our theoretical results, we measure the similarity between the target function $f^{*}$ and a minimizer $f$ of $\mathcal{L}$, under the augmentation $\mathcal{T_{H}}$. The proof of Theorems 4.2 and 4.3 shows that the affine transformation between $f$ and $f^{*}$ takes the form $f=WQf^{\ast}+b$, where $W$ is a whitening matrix, $b$ is the mean of the representations and $Q$ is an orthogonal matrix. Thus, in the case where $\text{cov}([f^{*}(x_{1}),\ldots,f^{*}(x_{n})])=I_{d}$ and the representations are centered, we expect $f=Qf^{*}$. We verify this in our next set of experiments by computing the average Procrustes distance between $\{f^{*}(x_{i})\}_{i=1}^{n}$ and $\{f(x_{i})\}_{i=1}^{n}$. It is defined as $\min_{Q\in\mathbb{R}^{d},Q^{\top}Q=I_{d}}\frac{1}{n}\|F-QF^{*}\|_{F}$ where $\|\cdot\|_{F}$ is the Frobenius norm, $F^{*}=[f^{*}(x_{1}),\ldots,f^{*}(x_{n})]$ and $F=[f(x_{1}),\ldots,f(x_{n})]$.
(a) Input-space augmentations
(b) Spectral Constrastive Loss
(c) VICReg
Figure 2: (a) Comparison of original and augmented MNIST images for different kernels. Notably, different function classes require different augmentations to achieve the same representations. (b)-(c) The average Procrustes distance between the learned representations and target/random representations during training given the augmentations $\mathcal{T}_{\mathcal{H}}$. We consider SCL, VICReg as defined in Section 3 as well as the original VICReg; we achieve the target representations up to equivalence for all losses.
SCL, Theoretical VICReg, and VICReg.
In Figures 2(b)-2(c), We optimize $\mathcal{L}_{\text{SCL}}$, $\mathcal{L}_{\text{VIC}}$ and the original VICReg loss (Bardes et al., 2022) under the augmentation $\mathcal{T}_{\mathcal{H}}$ with target representations obtained from a ResNet50 pretrained on ImageNet. We compute the target representations for 10,000 MNIST images (LeCun et al., 2010) and use the RBF kernel. We compare the Procrustes distance of the learned representations to the target representations, and to random representations with the same covariance. We observe that the representations learned with the constructed augmentation indeed achieve the target representations. It may seem surprising that the theoretical VICReg corresponds so closely to the original version, we explain this fact in the appendix. The final average Procrustes distances are 0.6 ± 0.1 for VICReg, 0.4 ± 0.1 for $\mathcal{L}_{\text{VIC}}$ and 0.004 ± 0.001 for $\mathcal{L}_{\text{SCL}}$.
Additional Experiments.
We repeat the experiments for a ViT-B/16 (Dosovitskiy et al., 2021) using CIFAR-10 (Krizhevsky, 2009) and Tiny ImageNet (Le and Yang, 2015). We optimize VICReg and measure the average procrustes distance after 300 epochs. The random baseline distance is 120.
Model MNIST CIFAR-10 Tiny ImageNet
ResNet50 0.3 ± 0.2 0.14 ± 0.04 0.2 ± 0.2
ViT-B/16 0.7 ± 0.1 0.7 ± 0.2 0.4 ± 0.2
7 Discussion
Our work is the first to “invert” the typical mathematical analysis of SSL, asking what the optimal augmentation for a desired target representation must be. We answer this question for two of the most popular non-contrastive losses (VICReg and Barlow Twins) and the Spectral Contrastive Loss — a theoretical proxy to SimCLR. Below, we discuss the main implications of our analysis.
The Role of Augmentations.
Augmentations are commonly thought of as providing different views of the same data. Our theoretical analysis provides a more nuanced look at this matter. Theorems 4.2 and 4.3 allow to interpret augmentations as projections in the feature space to a subspace, which guarantees the minimizer $f$ of the SSL loss to be equivalent a specific $f^{*}$, meaning that augmentations can act as projections in the feature space instead of generators of different views. Moreover, our results show that even one non-trivial augmentation is sufficient to learn good representations, and the distribution induced by this augmentation need not align with the original data distribution. This provides a first theoretical justification for recent empirical observations (Moutakanni et al., 2024).
The Impact of Architecture.
Figure 2(a) shows that for different architectures, the augmentations are recognizably different. This follows directly from our theory since the matrix $K$ differs depending on the learnable function class in our setup. Saunshi et al. (2022) note that architecture has an implicit bias on the representations, and that this affects downstream performance. Our work also sheds light on the architectural influence in SSL, but from a new perspective: For the same target representation, the function class heavily influences the optimal augmentations needed to learn these representations.
The choice of $f^{*}$.
Assuming a target $f^{*}$ is standard in most theoretical analyses of supervised learning, and also common in SSL theory (Simon et al., 2023; Zhai et al., 2024). By assuming an arbitrary target, we show that our results hold for any $f^{*}$. While we choose $f^{*}$ to be a pretrained model in our experiments, there is no general consensus on what constitutes a good representation. When there is a natural target, our theoretical framework allows investigating the interplay between augmentations and learned representations from a new angle . As a demonstrating example, suppose the data-generating process is governed by a statistical model, such as the spiked covariance model commonly studied in high-dimensional statistics (Wainwright, 2019): $x\sim\mathcal{N}_{m}(0,\Sigma);\Sigma=\nu\cdot\theta\theta^{\top}+I_{m}$ where $\nu>0$ and $\theta\in\mathbb{R}^{m}$ is the (unknown) signal direction hidden in the covariance $\Sigma$. Previous works on SSL would only be able to answer what representations $\hat{\theta}^{\top}x$ are learned from a given augmentation. In contrast, our theoretical results from Section 4 would give us analytical expressions for what the “correct” augmentation is, i.e. the one that projects onto the signal direction $\theta$.
Connection Between VICReg and SCL.
Our results show that both VICReg and SCL can achieve the same $f^{*}$ given the same augmentations. The connection between VICReg and SCL has been shown before for the limit $n\rightarrow\infty$ (Cabannes et al., 2023), and in the finite sample case, through replacing $c(Z)$ from the definition of VICReg with contrastive terms while keeping the non-contrastive variance term $v(Z)$ (Garrido et al., 2022). To the best of our knowledge, we are the first to show a theoretical connection between VICReg and SCL for finite data by regularizing only with contrastive terms, thereby providing strong support to the notion of duality between contrastive and non-contrastive learning (Garrido et al., 2022). Compared to previous works, we do not attempt to achieve functional equivalence between the losses, but instead show that the same augmentations can lead to the same $f^{*}$ when minimizing both losses, even if the optimization landscape is different.
Extension to Neural Networks.
The most significant limitation of our results is that, although neural networks can serve as the target $f^{*}$, the function class is assumed to be a kernel model, similar to previous theoretical studies on SSL. Our framework can extend, however, to neural networks as follows: Consider a parameterized function class $\Theta\phi_{\theta}(\cdot)$ and a target $f^{*}(x)=\Theta^{*}\phi_{\theta^{*}}(x)$. Taking the kernel $k_{\theta^{*}}(x,x^{\prime})=\phi_{\theta^{*}}(x)^{\top}\phi_{\theta^{*}}(x^{\prime})$, our theory provides augmentations such that $f^{*}$ uniquely optimizes the “frozen” neural network $\Theta\phi_{\theta^{*}}(\cdot)$, up to an affine transformation. Interestingly, $f^{*}$ will also be a global minimum over the “unfrozen” neural network $\Theta\phi_{\theta}(\cdot)$, as the proofs of Theorems 4.2, 4.3, 4.5 show that $f^{*}$ achieves the minimum possible value of the loss. The uniqueness of the minimum can only possibly be guaranteed in the infinite width limit, which is usually the case for neural networks, since otherwise $\phi_{\theta}$ changes during training. Nevertheless, it is encouraging that $f^{*}$ is a global minimizer also for finite-width neural networks, which motivates future research on the conditions under which $f^{*}$ or an equivalent function is achieved instead of other minima.
Acknowledgments and Disclosure of Funding
This work started as part of Shlomo Libo Feigin’s Master’s thesis at the Technical University of Munich. It was partially supported by the German Research Foundation (DFG) through the Research Grant GH 257/4-1, and by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research (BMBF). We thank Vladimir Cherkassky, Daniel He, Daniel Bin Schmid, Nimrod de la Vega and Maedeh Zarvandi for the interesting and helpful discussions on this work. We also wish to thank Nil Ayday, Alexandru Craciun and Pascal Esser for the useful comments on the paper.
Appendix A Definition of the VICReg Loss
Although in Section 3 we use the variance in the definition of $v$ as part of the VICReg loss, we can use a variety of other definitions for our results. For example, we can use the standard deviation similar to the original VICReg definition [Bardes et al., 2022]. The part of the proof where the definition of VICReg comes to play is Lemma C.2, where we use the fact that:
$v(Z)=0\iff\forall_{i\in[d]}\text{cov}(Z)_{i,i}=1$
Which is, of course, also true for:
$v(Z)=\frac{1}{d}\sum_{i=1}^{d}\left(1-\sqrt{[\text{cov}\left(Z\right)]_{i,i}}\right)^{2}$
Similar to the definition in Bardes et al. [2022], except we use the $L_{2}$ loss instead of the hinge loss. ,In fact, any loss function $l(\cdot)$ where $l(y)=0\iff y=0$ would work for our results.
In addition, the unnormalized cross-correlation can be used instead of the covariance (Theorem C.4).
Appendix B Proof of Statements in Section 3.2
B.1 $\mathcal{L}_{\text{SCL}}$ Can Diverge Without Regularization
The basic intuition behind this argument is that if we manage align the negative examples such that they have a close to 00 inner-product, then the norm can grow and the contibution of the term $-\frac{2}{n}{}\sum_{i=1}^{n}z_{i}^{\top}z^{\prime}_{i}$ will outweigh the contribution of the term $\frac{1}{n^{2}}\sum_{i\neq j}(z_{i}^{\top}z^{\prime}_{j})^{2}$.
More formally, to demonstrate that $\mathcal{L}_{\text{SCL}}$ can diverge to $\infty$- ∞ without any regularization, we look at the following setup:
Let $\{x_{i}\}_{i=1}^{n}\subseteq\mathcal{X}$ be the data divided into $b=\frac{n}{d}\in\mathbb{N}$ batches $B_{j}=\{x_{i}\}_{i=j\cdot n+1}^{(j+1)\cdot n}$, where $d$ is the dimension of the representation and $\mathcal{X}$ is assumed to be a compact metric space. Let $\mathcal{T}$ be a distribution of augmentations $T:\mathcal{X}\rightarrow\mathcal{X}$ with a finite support and let $\mathcal{F}$ be a universal function class, i.e. for every continuous function $g:\mathcal{X}\rightarrow\mathbb{R}^{d}$ in $\mathcal{X}\rightarrow\mathbb{R}^{d}$ and every $\epsilon>0$, there exists $f\in\mathcal{F}$ such that $\sup_{x\in\mathcal{X}}|f(x)-g(x)|_{2}<\epsilon$. Assume the augmentations do not overlap, i.e. $\forall_{T,T^{\prime}\in\mathcal{T}}\forall_{i\neq j}T(x_{i})\neq T^{\prime}(x_{j})$.
We define the unregularized loss as:
$\mathcal{L}_{\text{SCL-unreg}}(f)$ $=\sum_{j=1}^{b}\mathbb{E}_{Z_{j},Z_{j}^{\prime}}[L_{\text{SCL-unreg}}(Z_{j},Z_{j})]\quad\text{, where for a batch size $\nu$:}$
$L_{\text{SCL-unreg}}(Z,Z^{\prime})$ $=-\frac{2}{\nu}\sum_{i=1}^{\nu}z_{i}^{\top}z^{\prime}_{i}+\frac{1}{\nu^{2}}\sum_{i\neq j}(z_{i}^{\top}z^{\prime}_{j})^{2}$
and $Z_{j}$ corresponds to the $j$’s batch. Namely, for $\left(T_{i},T^{\prime}_{i}\right){\sim}\mathcal{T}\otimes\mathcal{T}$:
$Z_{j}$ $=[f(T_{j\cdot d+1}(x_{j\cdot d+1})),\ldots,f(T_{(j+1)\cdot d}(x_{(j+1)\cdot d}))]$
$Z_{j}^{\prime}$ $=[f(T_{j\cdot d+1}^{\prime}(x_{j\cdot d+1})),\ldots,f(T_{(j+1)\cdot d}^{\prime}(x_{(j+1)\cdot d}))]$
Claim.
In the setting above, $\inf_{f\in\mathcal{F}}\mathcal{L}_{\text{SCL-unreg}}(f)=-\infty$
Proof.
For a $\lambda>0$ Let $g(x)$ be a function such that $\forall_{T\in\text{Support}(\mathcal{T})}g(T(x_{i}))=\sqrt{\frac{\lambda}{2b}}\cdot e_{i\bmod d}$, where $e_{k}$ is the $k$’th standard basis vector. $g_{\lambda}(\cdot)$ is well defined on the finite set $S=\{T(x_{i}):i\in[n],T\in\text{Support}(\mathcal{T})\}$ since we assumed the augmentations do not overlap, i.e. $\forall_{T,T^{\prime}\in\mathcal{T}}\forall_{i\neq j}T(x_{i})\neq T^{\prime}(x_{j})$. The function $g_{\lambda}$ can be extended from the space $S$ to $\mathcal{X}$, such that $g_{\lambda}$ is continuous, via the Tietze extension theorem. Therefore, the universality of $\mathcal{F}$ implies that for every $\epsilon>0$ there exists an $f\in\mathcal{F}$ such that $\sup_{x\in\mathcal{X}}|f(x)-g(x)|_{2}<\epsilon$
Now, we can prove $\mathcal{L}_{\text{SCL-unreg}}(g_{\lambda})=-\lambda$:
$Z_{j}$ $=[g_{\lambda}(T_{j\cdot d+1}(x_{j\cdot d+1})),\ldots,g_{\lambda}(T_{(j+1)\cdot d}(x_{(j+1)\cdot d}))]$
$=[\sqrt{\frac{\lambda}{2b}}\cdot e_{1},\ldots\sqrt{\frac{\lambda}{2b}}\cdot e_{d}]=\sqrt{\frac{\lambda}{2b}}I_{d}$
and similarly $Z_{j}^{\prime}=\sqrt{\frac{\lambda}{2b}}I_{d}$. Therefore, $L_{\text{SCL-unreg}}(Z,Z^{\prime})=-\frac{2}{d}\sum_{i=1}^{d}e_{i}^{\top}e_{i}+\frac{1}{d^{2}}\sum_{i\neq j}e_{i}^{\top}e_{j}=-\frac{\lambda}{b}$ and $\mathcal{L}_{\text{SCL-unreg}}(g_{\lambda})=\sum_{j=1}^{b}\mathbb{E}_{Z_{j},Z_{j}^{\prime}}[L_{\text{SCL-unreg}}(Z_{j},Z_{j})]=-b\cdot\frac{\lambda}{b}=-\lambda$.
$\mathcal{L}_{\text{SCL-unreg}}(f)$ can be viewed as a continuous function from the set $\{f(T(x_{i})):T\in\text{Support}(\mathcal{T}),i\in[n]\}$ and therefore the universality of $\mathcal{F}$ implies that for every $\delta>0$ there is a $f\in\mathcal{F}$ such that $\mathcal{L}_{\text{SCL-unreg}}(f_{\lambda})\leq\mathcal{L}_{\text{SCL-unreg}}(g_{\lambda})+\delta$, in particular, we can fix $\delta=1$ and get that for every $\lambda>0$ there exists $f_{\lambda}\in\mathcal{F}$ such that $\mathcal{L}_{\text{SCL-unreg}}(f_{\lambda})\leq-\lambda+1$. This allows to prove our claim that $\inf_{f\in\mathcal{F}}\mathcal{L}_{\text{SCL-unreg}}(f)=-\infty$ by the following standard argument: Assume by contradiction $\inf_{f\in\mathcal{F}}\mathcal{L}_{\text{SCL-unreg}}(f)=M$ where $M\in\mathbb{R}$. For $\lambda=M-2$ we get $\exists_{f\in\mathcal{F}}\mathcal{L}_{\text{SCL-unreg}}(f)\leq M-1$ which contradicts the assumption $\inf_{f\in\mathcal{F}}\mathcal{L}_{\text{SCL-unreg}}(f)=M$ and therefore proving $\inf_{f\in\mathcal{F}}\mathcal{L}_{\text{SCL-unreg}}(f)=-\infty$. ∎
B.2 Lower Bound on $\mathcal{L}_{\text{SCL}}$
We begin by defining:
$L_{\text{SCL}}(Z,Z^{\prime})\coloneqq\|Z^{\top}Z^{\prime}-\text{diag}(Z^{\top}Z^{\prime})\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2n^{2}}\Big{(}\sum_{i=1}^{n}\|z_{i}\|^{4}+\|z_{i}^{\prime}\|^{4}\Big{)}$
It can easily be verified that $\mathcal{L_{\text{SCL}}}=\mathbb{E}_{Z,Z^{\prime}}\big{[}L_{\text{SCL}}(Z,Z^{\prime})\big{]}$. We will show that $\mathcal{L}_{\text{SCL}}\geq-d$ by showing that $L_{\text{SCL}}(Z,Z^{\prime})\geq-d$.
$L_{\text{SCL}}(Z,Z^{\prime})=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}-\text{diag}(Z^{\top}Z^{\prime})\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2n^{2}}\Big{(}\sum_{i=1}^{n}\|z_{i}\|^{4}+\|z_{i}^{\prime}\|^{4}\Big{)}$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{2}{n^{2}}\underbrace{\text{Tr}\big{(}(Z^{\top}Z^{\prime})^{\top}\text{diag}(Z^{\top}Z^{\prime})\big{)}}_{\|\text{diag}(Z^{\top}Z^{\prime})\|_{F}^{2}}+\frac{1}{n^{2}}\|\text{diag}(Z^{\top}Z^{\prime})\|_{F}^{2}$
$-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2n^{2}}\Big{(}\sum_{i=1}^{n}\|z_{i}\|^{4}+\|z_{i}^{\prime}\|^{4}\Big{)}$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2n^{2}}\Big{(}\sum_{i=1}^{n}\|z_{i}\|^{4}+\|z_{i}^{\prime}\|^{4}\Big{)}-\frac{1}{n^{2}}\|\text{diag}(Z^{\top}Z^{\prime})\|_{F}^{2}$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2n^{2}}\Big{(}\sum_{i=1}^{n}\|z_{i}\|^{4}-2\|z_{i}^{\top}z^{\prime}_{i}\|^{2}+\|z_{i}^{\prime}\|^{4}\Big{)}$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2n^{2}}\Big{(}\sum_{i=1}^{n}\|z_{i}z_{i}^{\top}-z_{i}^{\prime}z_{i}^{\prime\top}\|^{2}_{F}\Big{)}$
$\geq\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2n^{2}}\Big{(}\|\sum_{i=1}^{n}z_{i}z_{i}^{\top}-z_{i}^{\prime}z_{i}^{\prime\top}\|^{2}_{F}\Big{)}\quad$ (1)
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2}\Big{(}\|\frac{1}{n}ZZ^{\top}-\frac{1}{n}Z^{\prime}Z^{\prime\top}\|^{2}_{F}\Big{)}$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2}\Big{(}\|(\frac{1}{n}ZZ^{\top}-I_{d})-(\frac{1}{n}Z^{\prime}Z^{\prime\top}-I_{d})\|^{2}_{F}\Big{)}$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2}\|\frac{1}{n}ZZ^{\top}-I\|_{F}^{2}+\frac{1}{2}\|\frac{1}{n}Z^{\prime}Z^{\prime\top}-I_{d}\|_{F}^{2}$
$-\text{Tr}\Big{(}(\frac{1}{n}ZZ^{\top}-I_{d})^{\top}(\frac{1}{n}Z^{\prime}Z^{\prime\top}-I_{d})\Big{)}$
(1) is a consequence of the triangle inequality.
$\text{Tr}\Big{(}(\frac{1}{n}ZZ^{\top}-I)^{\top}(\frac{1}{n}Z^{\prime}Z^{\prime\top}-I)\Big{)}$
$=\frac{1}{n^{2}}\underbrace{\text{Tr}(ZZ^{\top}Z^{\prime}Z^{\prime\top})}_{=\text{Tr}(Z^{\prime\top}ZZ^{\top}Z^{\prime})}-\frac{1}{n}\text{Tr}(ZZ^{\top})-\frac{1}{n}\text{Tr}(Z^{\prime}Z^{\prime\top})+\text{Tr}(I_{d})$
$=\frac{1}{n^{2}}\text{Tr}\big{(}(Z^{\top}Z^{\prime})^{\top}(Z^{\top}Z^{\prime})\big{)}-\frac{1}{n}\text{Tr}(ZZ^{\top})-\frac{1}{n}\text{Tr}(Z^{\prime}Z^{\prime\top})+\text{Tr}(I_{d})$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{1}{n}\text{Tr}(ZZ^{\top})-\frac{1}{n}\text{Tr}(Z^{\prime}Z^{\prime\top})+d$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{1}{n}\big{(}\text{Tr}(ZZ^{\top})-2\text{Tr}(Z^{\top}Z^{\prime})+\text{Tr}(Z^{\prime}Z^{\prime\top})\big{)}+\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+d$
$=\frac{1}{n^{2}}\|Z^{\top}Z^{\prime}\|_{F}^{2}-\frac{1}{n}\|Z-Z^{\prime}\|_{F}^{2}+\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+d$
Combining everything together, we get:
$L_{\text{SCL}}(Z,Z^{\prime})\geq\frac{1}{2}\|\frac{1}{n}ZZ^{\top}-I\|_{F}^{2}+\frac{1}{2}\|\frac{1}{n}Z^{\prime}Z^{\prime\top}-I_{d}\|_{F}^{2}+\frac{1}{n}\|Z-Z^{\prime}\|_{F}^{2}-d\geq-d$ (2)
The inequality is tight if and only if $\frac{1}{n}ZZ^{\top}=I_{d}$ and $Z=Z^{\prime}$, because only then both (1) and the right side of (2) are equalities.
Appendix C Proofs of the Main Results
Additional Notation and Definitions.
In addition to the notation in Section 3, we denote $\phi_{i}\coloneqq\phi(x_{i})$, $\|\cdot\|_{\text{HS}}$ to be the Hilbert-Schmidt norm, and $O(n)$ to be the set of orthogonal $n\times n$ matrices. Moreover, we define $L_{\text{VIC}},L_{\text{BT}},L_{\text{SCL}}$ such that the loss functions in Section 3.2 are of the form $\mathcal{L}=\mathbb{E}_{Z,Z^{\prime}}\big{[}L(Z,Z^{\prime})\big{]}$, namely:
$L_{\text{VIC}}(Z,Z^{\prime})$ $\coloneqq\lambda s(Z,Z^{\prime})+\mu[v(Z)+v(Z^{\prime})]+\nu[c(Z)+c(Z^{\prime})]$
$L_{\text{BT}}(Z,Z^{\prime})$ $\coloneqq\sum_{i}^{n}(1-\mathcal{C}_{ii})^{2}+\lambda\sum_{i\neq j}(1-\mathcal{C}_{ij})^{2};\quad\mathcal{C}=\frac{1}{2n}(ZZ^{\prime\top}+Z^{\prime}Z^{\top})$
$L_{\text{SCL}}(Z,Z^{\prime})$ $\coloneqq\|Z^{\top}Z^{\prime}-\text{diag}(Z^{\top}Z^{\prime})\|_{F}^{2}-\frac{2}{n}\text{Tr}(Z^{\top}Z^{\prime})+\frac{1}{2n^{2}}\Big{(}\sum_{i=1}^{n}\|z_{i}\|^{4}+\|z_{i}^{\prime}\|^{4}\Big{)}$
C.1 Proof of Theorem 4.2 (VICReg)
To prove Theorem 4.2 we need a number of auxiliary results. The main part of the proof is in Proposition C.3, where we prove our main result except we assume $\text{cov}\left([{f}^{*}(\phi_{1}),\ldots,{f}^{*}(\phi_{n})]\right)=I_{d}$ instead of the covariance being full rank (Condition 3.2), which is easily relaxed afterwards.
First, we derive a representer-like theorem that guarantees any $f$ that minimizes VICReg with least norm to be of the form $f=A\Phi^{\top}$, provided a suitable augmentation.
Proposition C.1.
Let $\mathcal{L}$ be a joint embedding loss as described in Section 3. Let $\mathcal{T}$ be a distribution of augmentations such that $\forall_{T\in\text{Support}(\mathcal{T})}T\Phi\in\{\Phi AK:A\in\mathbb{R}^{n\times n}\}$. Suppose $f\in\arg\min\{\|f\|_{HS}:\mathcal{L}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},f)=0\}$. Then, $f=A\Phi^{\top}$ for some $A\in\mathbb{R}^{d\times n}$.
Proof.
Let $\mathcal{T}$ be a distribution of augmentions s.t. $\forall_{T\in\text{Support}(\mathcal{T})}T\Phi\in\{\Phi AK:A\in\mathbb{R}^{n\times n}\}$ and let $f\in\mathbb{R}^{d}\otimes\mathcal{H}$ be a solution to ($\mathcal{L}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},f)=0$ Define $S_{\parallel}\coloneqq\{A\Phi^{\top}:A\in\mathbb{R}^{d\times n}\}$. We decompose $f=f_{\parallel}+f_{\perp}$ where $f_{\parallel}\in S_{\parallel}$ and $<f_{\perp},\hat{f}>_{\text{HS}}=0$ for every $\hat{f}\in S_{\parallel}$. Since $S_{\perp}$ is finite-dimensional, such a decomposition exists via the Gram-Schmidt process.
$\mathcal{L}$ is a function of $\{f(T(\phi_{i}):T\in\text{Support}(\mathcal{T}),i\in[n]\}$, and since we assume $\forall_{T\in\text{Support}(\mathcal{T})}T\Phi\in\{\Phi AK:A\in\mathbb{R}^{n\times n}\}$, we get that $\mathcal{L}$ is a function of $\{f(\phi_{i})\}_{i=1}^{n}$.
We first prove that $\forall_{i\in[n]}f_{\perp}(\phi_{i})=0$. Let $A\in\mathbb{R}^{d\times n}$ be an arbitrary matrix.
$0=<f_{\perp},A\Phi^{\top}>_{\text{HS}}=\text{Tr}(f_{\perp}^{\top}A\Phi^{\top})=\text{Tr}(A\Phi^{\top}f_{\perp}^{\top})=\sum_{i=1}^{n}<Ae_{i},\Phi^{\top}f_{\perp}^{\top}e_{i}>$
Where $\Phi^{\top}f_{\perp}^{\top}$ is a $n\times n$ matrix. Since $A$ is arbitrary it means that $\Phi^{\top}f_{\perp}^{\top}=0$ and hence $(\Phi^{\top}f_{\perp}^{\top})^{\top}=f_{\perp}\Phi=0$ Therefore: $\forall_{i\in[n]}f_{\perp}(\phi_{i})=0$. We get that for every $i\in[n]$:
$f(\phi_{i})$ $=f_{\parallel}(\phi_{i})$
$\|f\|_{\text{HS}}^{2}$ $=\|f_{\parallel}\|_{\text{HS}}^{2}+\|f_{\perp}\|_{\text{HS}}^{2}$
And since $f$ is a min-norm solution of ($\mathcal{L}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},f)=0$ it must satisfy $f=f_{\parallel}\in S_{\parallel}$.
We continue by deriving two conditions that are equivalent to obtaining zero loss in VICReg.
Lemma C.2.
$L_{\text{VIC}}(Z,Z^{\prime})=0$ if and only if $Z=Z^{\prime}$ and $\text{cov}(Z)=n^{-1}ZHH^{\top}Z^{\top}=I$.
Proof.
$L_{\text{VIC}}(Z,Z^{\prime})=0$ if and only if:
$s(Z,Z^{\prime})$ $=\frac{1}{n}\sum_{i=1}^{n}\left\|z_{i}-z_{i}^{\prime}\right\|_{2}^{2}=0$
$v(Z)$ $=\frac{1}{d}\sum_{i=1}^{d}\left(1-[\text{cov}\left(Z\right)]_{i,i}\right)^{2}=0$
$c(Z)$ $=\frac{1}{d}\sum_{i\neq j}^{d}[\text{cov}(Z)]_{i,j}^{2}=0$
Which is satisfied if and only if $Z=Z^{\prime}$ and $\text{cov}(Z)=I$.
Next, we show a statement that is already very close to Theorem 4.2. The only difference here is that we still assume the target representations to have identity covariance. This simplifies the proof, but it will be easy to relax.
Proposition C.3.
Suppose Condition 3.3 is satisfied. Let $f^{*}\in\mathbb{R}^{d}\otimes\mathcal{H}$ be of the form $f^{*}=C\Phi^{\top},C\in\mathbb{R}^{d\times n}$ s.t. $\text{cov}\left([f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]\right)=`I_{d}$. Then, if $f$ minimizes $\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T}_{\mathcal{H}}(C),f\right)$ over $\mathbb{R}^{d}\otimes\mathcal{H}$ with least Hilbert-Schmidt-norm, there exists $Q\in O(d)$ s.t. $f=Qf^{*}$.
Proof.
Our goal in this proof is to show that the set of least norm minimizers of $\mathcal{L}_{\text{VIC}}$, which will be denoted by $\mathcal{F^{*}}$, is the same as $\{Qf^{*}:Q\in O(d)\}$. The general structure of the proof is as follows: We first prove that any $Qf^{*}$ indeed achieves minimum loss (Claim 1) and then prove that any least norm minimizer must be of this form (Claim 2) — which is the crux of the proof. Claims 1 and 2 will not be sufficient by themselves to prove the equivalence between $\{Qf^{*}:Q\in O(d)\}$ and $\mathcal{F^{*}}$, since it could still be that there is a $Qf^{*}$ that minimizes the loss (which we know by Claim 1) but is not a least norm solution. We prove that this cannot be the case in Claim 3 and thereby prove the proposition.
We first show that $\mathcal{T}_{\mathcal{H}}(C)$ is well defined by proving that $CKC^{\top}$ has full rank. This follows from the fact that if $K$ is a positive definite symmetric matrix (Condition 3.3), and if $K=LL^{\top}$ is the Cholesky decomposition of $K$, $CKC^{\top}=(CL)(CL)^{\top}$ and $rank(CKC^{\top})=rank(CL)=rank(C)$. The fact that $rank(C)=d$ follows our assumption that: $\text{cov}\left([f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]\right)=\text{cov}(C\Phi^{\top}\Phi)=\frac{1}{n}(C\Phi^{\top}\Phi H)(C\Phi^{\top}\Phi H)^{\top}=I_{d}$ and because of the submultiplicativity of the matrix rank $rank(C)\geq rank(I_{d})=d$. Of course $rank(C)\leq d$ because $C$ is a $d\times n$ matrix. To summarize, we get $rank(CKC^{\top})=rank(C)=d$.
For simplicity of notation we denote $\mathcal{T}=\mathcal{T}_{\mathcal{H}}(C)$. Define:
$\mathcal{F}^{\ast}\coloneqq\arg\min_{f\in\mathbb{R}^{d}\otimes\mathcal{H}}\{\|f\|_{\text{HS}}:\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$
We will prove that =$\mathcal{F}^{*}=\{QC\Phi^{\top}:Q\in O(d)\}$. Since $\mathcal{L}_{\text{VIC}}$ is non-negative, the property $\mathcal{L}_{\text{VIC}}=0$ also implies that $\mathcal{F}^{\ast}$ is the set of minimizers of $\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)$ with least Hilbert-Schmidt-norm.
The fact that $\mathcal{L}_{\text{VIC}}$ is non-negative also implies that $\mathcal{L}_{\text{VIC}}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},f)=\mathbb{E}_{Z,Z^{\prime}}\big{[}L_{\text{VIC}}(Z,Z^{\prime})\big{]}=0$ if and only if every term inside the expectation is exactly 00. i.e. for every $(T_{1},T^{\prime}_{1},\ldots,T_{n},T^{\prime}_{n})$ in the support of $(\mathcal{T})^{\otimes 2n}$
$L_{\text{VIC}}([f(T_{1}(\phi_{1})),\ldots,f(T_{n}(\phi_{n}))],[f(T^{\prime}_{1}(\phi_{1})),\ldots,f(T^{\prime}_{n}(\phi_{n}))])=0$ (3)
Claim 1.
We prove that $\{QC\Phi^{\top}:Q\in O(d)\}\subseteq\{f:\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$.
Let $f=QC\Phi^{\top}$ where $Q\in O(d)$. We prove that for every $T,T^{\prime}\in\text{Support}(\mathcal{T})$ and for every $i\in[n]$ it holds that $f(T(\phi_{i}))=f(T^{\prime}(\phi_{i}))$. Since $\text{Support}(\mathcal{T})=\{I_{\mathcal{H}},\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}\}$, it is enough to check that the following holds for every $i\in[n]$:
$QC\Phi^{\top}I_{\mathcal{H}}\phi_{i}=QC\Phi^{\top}(\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T})\phi_{i}$
Which indeed holds since:
$QC\Phi^{\top}\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}\phi_{i}=Q\underbrace{CKC^{T}\left(CKC^{T}\right)^{-1}}_{I}C\Phi^{T}\phi_{i}=QC\Phi^{T}\phi_{i}$
Therefore, (3) is exactly:
$L_{\text{VIC}}(QC\Phi^{T}\Phi,QC\Phi^{T}\Phi)=0$
$L_{\text{VIC}}(QCK,QCK)=0$
Which according to Lemma (C.2) holds if $\text{cov}(QCK)=I$, which is true since we assumed that $\text{cov}(CK)=I$ and therefore indeed $L_{\text{VIC}}(QCK,QCK)=0$, (3) holds and Claim 1 is proven.
Claim 2.
We prove that $\mathcal{F}^{\ast}\subseteq\{QC\Phi^{\top}:Q\in O(d)\}$.
Let $f\in\mathcal{F}^{\ast}$. The structure of the proof will be as follows: First, we prove that $f$ is of the form $f=A\Phi^{\top}$ via the version of the representer theorem we proved (Proposition C.1). Then, we use Lemma C.2 which states the conditions under which $L(Z,Z^{\prime})=0$ and substituting different possible augmentations combinations that could occur under $\mathcal{T}$ into $Z,Z^{\prime}$ and get equations (4) and (5) (we use the fact that no matter the combination, $L(Z,Z^{\prime})=0$, which is formalized by eq. 3). Combining (1), (5) and the assumption that the covariance is identity will force $A$ to be of the form $A=QC$, which proves the claim.
To start, to prove $f=A\Phi^{\top}$ we would like to use Proposition C.1, which requires $\forall_{T\in\text{Support}(\mathcal{T})}T\Phi\in\{\Phi\bar{A}K:\bar{A}\in\mathbb{R}^{n\times n}\}$. For $T=\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}$ the condition is satisfied since by definition $K=\Phi^{\top}\Phi$. For $T=I_{\mathcal{H}}$ the condition is satisfied for $\bar{A}=K^{-1}$. Therefore, according to Proposition C.1 $f=A\Phi^{\top}$.
We now use (3). According to Lemma C.2 for every $i\in[n]$ and $T,T^{\prime}\in\text{Support}(\mathcal{T})$:
$A\Phi^{\top}T_{i}(\phi_{i})=A\Phi^{\top}T_{i}^{\prime}(\phi_{i})$
Taking $T_{i}=I_{\mathcal{H}}$ and $T_{i}^{\prime}=\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}$ results in the following:
$A\Phi^{\top}I_{\mathcal{H}}(\phi_{i})=A\Phi^{\top}\Phi C^{T}\left(CKC^{T}\right)^{-1}C\Phi^{T}\phi_{i}$
Which holds for every $i\in[n]$, meaning:
$AK=AKC^{T}\left(CKC^{T}\right)^{-1}CK$ (4)
We use Lemma C.2 again for (3) to obtain that for $T_{1}=\ldots=T_{n}=T^{\prime}_{1}=\ldots=T^{\prime}_{n}=I_{\mathcal{H}}$
$\text{cov}(A\Phi^{\top}T_{1}(\phi_{1}),\ldots,A\Phi^{\top}T_{n}(\phi_{n}))=\text{cov}(A\Phi^{\top}\Phi)=\text{cov}(AK)=I$
$n^{-1}(AKH)(AKH)^{\top}=I$
$n^{-1}AKHH^{\top}KA^{\top}=I$ (5)
In addition, we assumed that $\text{cov}(CK)=I$, Therefore:
$n^{-1}CKHH^{\top}KC^{\top}=I$ (6)
We substitute (4) into (5) and get:
$n^{-1}AKC^{T}\left(CKC^{T}\right)^{-1}CKHH^{\top}KC^{\top}\left(CKC^{T}\right)^{-1}CKA^{\top}=I$
We substitute (6) into the above equation and get:
$AKC^{T}\left(CKC^{T}\right)^{-1}\left(CKC^{T}\right)^{-1}CKA^{\top}=I$ (7)
We denote $Q=AKC^{T}\left(CKC^{T}\right)^{-1}$, the above equation means that $Q\in O(d)$. Substituting $Q$ back to (4) we get:
$AK=QCK$
$A=QC$
Therefore, $f\in\{QC\Phi^{\top}:Q\in O(d)\}$ and Claim 2 is proven.
Claim 3.
$\{QC\Phi^{\top}:Q\in O(d)\}\subseteq\mathcal{F}^{\ast}$.
We assume by way of contradiction that there exists $f=\tilde{Q}C\Phi^{\top}\in\{QC\Phi^{\top}:Q\in O(d)\}$ such that $f\notin\mathcal{F}^{*}$. We know via Claim 1 that $\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},\tilde{f}\right)=0$, therefore the only way $f\notin\mathcal{F}^{*}$ is by not being a least-norm minimizer, i.e. there must be a $f^{\prime}\in\mathcal{F}^{\ast}$ such that $\|\ f^{\prime}\|_{\text{HS}}<\|f\|_{\text{HS}}$. Applying Claim 2 we know that $f^{\prime}=\tilde{Q}^{\prime}C\Phi^{\top}$ for $\tilde{Q}^{\prime}\in O(d)$. We get that $\|\tilde{Q}C\Phi^{\top}\|_{\text{HS}}<\|\tilde{Q}^{\prime}C\Phi^{\top}\|_{\text{HS}}$ for two matrices $\tilde{Q},\tilde{Q}^{\prime}\in O(d)$. Which is a contradiction since for every $Q\in O(d)$:
$\|QC\Phi^{\top}\|_{\text{HS}}^{2}=Tr((QC\Phi^{\top})^{\top}(QC\Phi^{\top}))=Tr((Q\Phi^{\top})^{\top}Q^{\top}Q(C\Phi^{\top}))$
$=Tr((C\Phi^{\top})^{\top}(C\Phi^{\top}))=\|C\Phi^{\top}\|_{\text{HS}}^{2}$
Therefore, we have proven Claim 3. Combining Claims 2 and 3 we get:
$\{QC\Phi^{\top}:Q\in O(d)\}=\mathcal{F}^{\ast}$
With Proposition C.3 in hand, we prove Theorem 4.2; this time, we do not assume a whitened covariance:
See 4.2
Proof.
[…]
C.2 Proof of Theorem 4.3 (Spectral Contrastive Loss)
The structure of our proof is as follows: First, we define a VICReg loss $\mathcal{L}_{\text{VIC-corr}}$ that uses the unnormalized cross-correlation instead of the covariance. We will verify that an equivalent version of Theorem 4.2 holds for $\mathcal{L}_{\text{VIC-corr}}$ with the same augmentations $\mathcal{T_{H}}$ as in Definition 4.1. Then, we will use the inequality derived in Section B.2 to show that $L(Z,Z^{\prime})_{\text{VIC-corr}}=0\iff{L}_{\text{SCL}}(Z,Z^{\prime})=-d$.
We start by defining:
$L_{\text{VIC-corr}}(Z,Z^{\prime})=\lambda s(Z,Z^{\prime})+\mu[$ $v_{\text{corr}}(Z)+v_{\text{corr}}(Z^{\prime})]+\nu[c_{\text{corr}}(Z)+c_{\text{corr}}(Z^{\prime})]\text{, where}$
$s(Z,Z^{\prime})$ $=\frac{1}{n}\sum_{i=1}^{n}\|z_{i}-z_{i}^{\prime}\|_{2}^{2},$
$v_{\text{corr}}(Z)$ $=\frac{1}{d}\sum_{i=1}^{d}\left(1-[\frac{1}{n}ZZ^{\top}]_{i,i}\right)^{2},$
$c_{\text{corr}}(Z)$ $=\frac{1}{d}\sum_{i\neq j}^{d}[\frac{1}{n}ZZ^{\top}]_{i,j}^{2}$
and $\lambda,\mu,\nu>0$ are hyper-parameters. We then define $\mathcal{L}_{\text{VIC-corr}}\Bigl{(}\{x_{i}\}_{i=1}^{n},\mathcal{T},f\Bigl{)}=\mathbb{E}_{Z,Z^{\prime}}\Bigl{[}L_{\text{VIC-corr}}(Z,Z^{\prime})\Bigr{]}$, $z_{i}=f\Bigl{(}T_{i}(x_{i})\Bigl{)},z_{i}^{\prime}=f\Bigl{(}T^{\prime}_{i}(x_{i})\Bigl{)}$, where $\left(T_{i},T^{\prime}_{i}\right){\sim}\mathcal{T}\otimes\mathcal{T}$.
An equivalent version of Theorem 4.2 holds for $\mathcal{L}_{\text{VIC-corr}}$, which we state below as Theorem C.4. Since we use Theorem C.4 as an auxiliary result for proving the optimality of augmentation for the Spectral Contrastive Loss, we formulate a stronger statement, where the loss of the minimizer is also guaranteed to be 00. However, as we will see, this does not significantly change the proof.
Theorem C.4 (Optimality of Augmentations for VICReg-corr).
[…]
The proof of this theorem is very similar to the proof of Theorem 4.2, and we will just note the exact differences.
The main difference between the proofs of Theorems 4.2 and C.4 lies in Lemma C.2, which is the only place the definition of $v(Z)$ and $c(Z)$ come into play. In the proof of Theorem C.4, Lemma C.2 is replaced with the following lemma:
Lemma C.5.
$L_{\text{VIC-corr}}(Z,Z^{\prime})=0$ if and only if $Z=Z^{\prime}$ and $\frac{1}{n}ZZ^{\top}=I$.
In the proof of Theorem 4.2, Lemma C.2 is used only in the proof of Proposition C.3, which in the proof of Theorem C.4 is replaced with the following proposition. Here, the covariance condition is replaced with a condition on the correlation, and the statement is stronger:
Proposition C.6.
Suppose Condition 3.3 is satisfied. Let $f^{*}\in\mathbb{R}^{d}\otimes\mathcal{H}$ be of the form $f^{*}=C\Phi^{\top},C\in\mathbb{R}^{d\times n}$ s.t. $\frac{1}{n}[f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})][f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]^{\top}=I_{d}$. Then, if $f$ minimizes $\mathcal{L}_{\text{VIC-corr}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T}_{\mathcal{H}}(C),f\right)$ with least Hilbert-Schmidt-norm over $\mathbb{R}^{d}\otimes\mathcal{H}$, $f$ satisfies $\mathcal{L}_{\text{VIC-corr}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T}_{\mathcal{H}}(C),f\right)=0$ and there exists $Q\in O(d)$ such that $f=Qf^{*}$.
Remark C.7.
The statement “$\mathcal{L}_{\text{VIC-corr}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T}_{\mathcal{H}}(C),f\right)=0$ and there exists $Q\in O(d)$ such that $f=Qf^{*}$” might seem stronger than the statement in Proposition C.3, which only guarantees there exists $Q\in O(d)$ such that $f=Qf^{*}$. However, the proof of Proposition C.3 showed that:
$\{QC\Phi^{\top}:Q\in O(d)\}=\mathcal{F}^{\ast},\quad\mathcal{F}^{\ast}\coloneqq\arg\min_{f\in\mathbb{R}^{d}\otimes\mathcal{H}}\{\|f\|_{\text{HS}}:\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$
and therefore because $\mathcal{F}^{*}$ is not empty, the proof also shows that if $f$ is a least norm minimizer then $\mathcal{L}_{\text{VIC}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$. Therefore, the stronger statement does not change the rest of the proof.
We proceed with detailing the changes that are caused by using Lemma C.5 instead of Lemma C.2. Namely, we replace eq. (5) with:
$n^{-1}AKKA^{\top}=I$ (8)
In addition, we assumed that $\frac{1}{n}[f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})][f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]^{\top}=I_{d}$ instead of $\text{cov}(CK)=I$, therefore insread of (6) we get:
$n^{-1}CKKC^{\top}=I$ (9)
We substitute (4) into (8) and get:
$n^{-1}AKC^{T}\left(CKC^{T}\right)^{-1}CKKC^{\top}\left(CKC^{T}\right)^{-1}CKA^{\top}=I$
We substitute (9) into the above equation and get:
$AKC^{T}\left(CKC^{T}\right)^{-1}\left(CKC^{T}\right)^{-1}CKA^{\top}=I$
This is exactly eq. (7), therefore the rest of the proof of Proposition C.6 goes exactly as Proposition C.3.
We now finish the proof of Theorem C.4.
Proof.
The only difference between Theorem C.4 and Proposition C.6 is that now instead of assuming $f^{*}=C\Phi^{\top},C\in\mathbb{R}^{d\times n}$ s.t. $\frac{1}{n}[f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})][f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]^{\top}=I_{d}$ we just assume $\text{cov}\left([f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]\right)$ it is full rank. The covariance being full rank implies $\frac{1}{n}[f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})][f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]^{\top}$ is full rank:
$rank(\frac{1}{n}[f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})][f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]^{\top})$
$=rank([f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})])$
$\geq rank([f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]H_{n})$
$=rank(\text{cov}\left([f^{*}(\phi_{1}),\ldots,f^{*}(\phi_{n})]\right)$
[…]
We thereby proved the optimality of augmentations for $\mathcal{L}_{\text{VIC}}$ and we proceed by connecting $\mathcal{L}_{\text{VIC-corr}}$ to $\mathcal{L}_{\text{SCL}}$.
Lemma C.8.
$\mathcal{L}_{\text{SCL}}\Bigl{(}\{(\phi_{i})\}_{i=1}^{n},\;\mathcal{T},\;f\Bigr{)}=-d\iff\mathcal{L}_{\text{VIC-corr}}\Bigl{(}\{(\phi_{i})\}_{i=1}^{n},\;\mathcal{T},\;f\Bigr{)}=0$
Proof.
The necessary and sufficient conditions for $L_{\text{VIC}}(Z,Z^{\prime})=0$ according to Proposition C.5 and the necessary and sufficient conditions on the tightness of the bound in Section B.2, which implies $L_{\text{SCL}}(Z,Z^{\prime})=-d\iff Z=Z^{\prime}\text{ and }\frac{1}{n}ZZ^{\top}=I\iff L_{\text{VIC-corr}}(Z,Z^{\prime})=0$. Since 00 and $-d$ are the lowest possible values the respective losses can take, we get a series of equivalent statements:
$\mathcal{L}_{\text{SCL}}\Bigl{(}\{(\phi_{i})\}_{i=1}^{n},\;\mathcal{T},\;f\Bigr{)}=0$
$\iff\forall_{(T_{1},T^{\prime}_{1},\ldots,T_{n},T^{\prime}_{n})\in\text{Support}(\mathcal{T}^{\otimes 2n})}L_{\text{SCL}}(Z,Z^{\prime})=-d$
$\iff\forall_{(T_{1},T^{\prime}_{1},\ldots,T_{n},T^{\prime}_{n})\in\text{Support}(\mathcal{T}^{\otimes 2n})}L_{\text{VIC-corr}}(Z,Z^{\prime})=0$
$\iff\mathcal{L}_{\text{VIC-corr}}\Bigl{(}\{(\phi_{i})\}_{i=1}^{n},\;\mathcal{T},\;f\Bigr{)}=0$
Where: $Z=[f(T_{1}(\phi_{1})),\ldots,f(T_{n}(\phi_{n}))]$ and $Z^{\prime}=[f(T^{\prime}_{1}(\phi_{1})),\ldots,f(T^{\prime}_{n}(\phi_{n}))]$. ∎
Now, we proceed by proving Theorem 4.3: See 4.3
Proof.
Let $f$ be a least norm minimizer of $\mathcal{L}_{\text{SCL}}((\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(C),\;f)$. We begin by arguing that $f$ is a least norm minimizer of $\mathcal{L}_{\text{VIC-corr}}(\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(C),\;f)$:
Let $f^{\prime}$ be a least norm minimizer of $\mathcal{L}_{\text{VIC-corr}}(\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(C),\;f)$, we know from Theorem C.4 that $\mathcal{L}_{\text{VIC-corr}}(\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(C),\;f^{\prime})=0$ and therefore from Lemma C.8 we know $\mathcal{L}_{\text{SCL}}(\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(C),\;f^{\prime})=-d$ and therefore $\mathcal{L}_{\text{SCL}}(\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(C),\;f)=-d$, since otherwise $f$ would not minimize $\mathcal{L}_{\text{SCL}}$. We use Lemma C.8 again in the other direction and get that $\mathcal{L}_{\text{VIC-corr}}(\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(C),\;f)=0$, therefore $f$ must be a minimizer of $\mathcal{L}_{\text{VIC-corr}}$. Moreover, $f$ must be a least norm minimizer of $\mathcal{L}_{\text{VIC-corr}}$, since any minimizer with a smaller norm would also minimize $\mathcal{L}_{\text{SCL}}$ (again, by Lemma C.8), contradicting the assumtion that $f$ is the least norm minimizer of $\mathcal{L}_{\text{SCL}}$.
[…]
C.3 Proof of Theorem 4.5 (Barlow Twins)
We structure the proof by first proving a set of auxiliary results and then proving Theorem 4.5.
Lemma C.9.
Let $T:\mathcal{H}\rightarrow\mathcal{H}$, and $\mathcal{T}$ a distribution of that yields $T$ with probability $0.5$ and $I_{\mathcal{H}}$ with probability $0.5$. Then, $\mathcal{L}_{BT}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{\top})=0$ if and only if $(2n)^{-1}A\Phi^{\top}(T\Phi\Phi^{\top}+\Phi\Phi^{\top}T^{\top})\Phi A^{\top}=I_{d}$.
Proof.
Since $\mathcal{L}_{\text{BT}}$ is non-negative, $\mathcal{L}_{\text{BT}}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{\top})=0$ if and only if for every $(T_{1},T^{\prime}_{1},\ldots,T_{n},T^{\prime}_{n})$ in the support of $(\mathcal{T}\otimes\mathcal{T}\mid T\neq T^{\prime})^{\otimes n}$:
$L_{\text{BT}}([A\Phi^{\top}T_{1}(\phi_{1}),\ldots,A\Phi^{\top}T_{n}(\phi_{n})],[A\Phi^{\top}T^{\prime}_{1}(\phi_{1}),\ldots,A\Phi^{\top}T^{\prime}_{n}(\phi_{n})])=0$
Recall the definition of $L_{\text{BT}}$:
$L_{\text{BT}}(Z,Z^{\prime})$ $=\sum_{i}^{n}(1-\mathcal{C}_{ii})^{2}+\lambda\sum_{i}^{n}(1-\mathcal{C}_{ii})^{2}$
Since $\lambda>0$, we get that $L_{\text{BT}}(Z,Z^{\prime})=0$ if and only if $\mathcal{C}\coloneqq\frac{1}{2n}(ZZ^{\prime\top}+Z^{\prime}Z^{\top})=I$. Plugging $ZZ^{\prime\top}=\sum_{i=1}^{n}(A\Phi^{\top}T_{i}(\phi_{1}))(A\Phi^{\top}T_{i}^{\prime}(\phi_{1}))^{\top}$:
$\frac{1}{2n}\sum_{i=1}^{n}(A\Phi^{\top}T_{i}(\phi_{i}))(A\Phi^{\top}T_{i}^{\prime}(\phi_{i}))^{\top}+(A\Phi^{\top}T^{\prime}_{i}(\phi_{i}))(A\Phi^{\top}T_{i}(\phi_{i}))^{\top}=I$ (10)
Since $\mathcal{T}$ has two augmentations in its support and we draw the augmentations conditioned on $T_{i}\neq T^{\prime}_{i}$, every summand of (10) will be of the following form:
$(A\Phi^{\top}T(\phi_{i}))(A\Phi^{\top}I_{\mathcal{H}}(\phi_{i}))^{\top}+(A\Phi^{\top}I_{\mathcal{H}}(\phi_{i}))(A\Phi^{\top}T(\phi_{i}))^{\top}$
Hence, we write (10) in the following way:
$\frac{1}{2n}\sum_{i=1}^{n}(A\Phi^{\top}T(\phi_{i}))(A\Phi^{\top}I_{\mathcal{H}}(\phi_{i}))^{\top}+(A\Phi^{\top}I_{\mathcal{H}}(\phi_{i}))(A\Phi^{\top}T(\phi_{i}))^{\top}$ $=I$
$\frac{1}{2n}\sum_{i=1}^{n}(A\Phi^{\top}T\phi_{i})(A\Phi^{\top}\phi_{i})^{\top}+(A\Phi^{\top}\phi_{i})(A\Phi^{\top}T\phi_{i})^{\top}$ $=I$
$\frac{1}{2n}\sum_{i=1}^{n}A\Phi^{\top}T\phi_{i}\phi_{i}^{\top}\Phi A^{\top}+A\Phi^{\top}\phi_{i}\phi_{i}^{\top}T^{\top}\Phi A^{\top}$ $=I$
$\frac{1}{2n}A\Phi^{\top}(T(\sum_{i=1}^{n}\phi_{i}\phi_{i}^{\top})+(\sum_{i=1}^{n}\phi_{i}\phi_{i}^{\top})T^{\top})\Phi A^{\top}$ $=I$
$\frac{1}{2n}A\Phi^{\top}(T\Phi\Phi^{\top}+\Phi\Phi^{\top}T^{\top})\Phi A^{\top}$ $=I$
Lemma C.10.
[See also Simon et al., 2023, Proposition 4.3] Let $\Gamma\in\mathbb{R}^{d\times d}$ be a symmetric positive definite matrix. Let $\Gamma=USU^{\top}$ its spectral decomposition. The set of solutions of $W\Gamma W^{\top}=I$ is exactly $W=QS^{-\frac{1}{2}}U^{\top}$ where $Q\in O(d)$.
Proof.
It is easy to verify that $W=QS^{-\frac{1}{2}}U^{\top}$ where $Q\in O(d)$ indeed satisfies $W\Gamma W^{\top}=I$. We would like to prove that any $W$ such that $W\Gamma W^{\top}=I$ is of that form. Denote $\Gamma^{\frac{1}{2}}=US^{\frac{1}{2}}U^{\top}$.
$(W\Gamma^{\frac{1}{2}})(W\Gamma^{\frac{1}{2}})^{\top}=I$
Hence, $W\Gamma^{\frac{1}{2}}\in O(d)$. Finally:
$W=W\Gamma^{\frac{1}{2}}(\Gamma^{\frac{1}{2}})^{-1}=W\Gamma^{\frac{1}{2}}US^{-\frac{1}{2}}U^{\top}$
Where $W\Gamma^{\frac{1}{2}}U\in O(d)$.
Lemma C.11.
Let $C$ be a rank $d$ matrix and $K$ be a positive definite matrix. The set of solutions $A$ to $AKC^{\top}(CKC^{\top})^{-2}CKA^{\top}=I_{d}$ is $\{QC:Q\in O(d)\}$
Proof.
Recall Equation (12):
$AKC^{\top}(CKC^{\top})^{-2}CKA^{\top}=I_{d}$
We denote $\Gamma=(CKC^{\top})^{-2}$. First, we show that the least Frobenius norm solutions of $W\Gamma W^{\top}$ are $W=QCKC^{\top}$.
$CKC^{\top}$ is a positive definite and symmetric, and $rank(CKC^{\top})=d$. This follows from the fact that if $K$ is a positive definite symmetric matrix, and if $K=LL^{\top}$ is the Cholesky decomposition of $K$, $CKC^{\top}=(CL)(CL)^{\top}$ and $rank(CKC^{\top})=rank(CL)=rank(C)=d$.
Let $V\Sigma V^{\top}$ be the spectral decomposition of $CKC^{\top}$.
$(CKC^{\top})^{-2}=V\Sigma^{-2}V^{\top}$
Therefore, according to Lemma C.10, the least Frobenius norm solutions of $W\Gamma W^{\top}$ are exactly $W=U\Sigma V^{\top}$.
$W=U\Sigma V^{\top}=UI_{d}\Sigma V^{\top}=UV^{\top}V\Sigma V^{\top}=(UV^{\top})CKC^{\top}.$
Where $UV^{\top}\in O(d)$. Consequently, all solutions to (12) are of the form $A=QC$ where $Q\in O(d)$.
See 4.5
Proof.
We prove a slightly stronger result, namely, instead of proving the equivalence of $f$ and $f^{*}$ up to an affine transformation, we prove an equivalence up to an orthogonal transformation, meaning there exists $Q\in O(d)$ s.t. $f=Qf^{*}$.
For simplicity of notation, we denote $\mathcal{T}=\mathcal{T}_{\mathcal{H}}^{\text{BT}}(C)$. Define:
$\mathcal{F}^{\ast}\coloneqq\arg\min_{f\in\mathbb{R}^{d}\otimes\mathcal{H}}\{\|f\|_{\text{HS}}:\mathcal{L}_{\text{BT}}\left(\left\{\phi_{i}\right\}_{i=1}^{n},\mathcal{T},f\right)=0\}$
We will prove that =$\mathcal{F}^{*}=\{QC\Phi^{\top}:Q\in O(d)\}$.
We begin by rewriting the kernel Gram matrix $K$ and the Lyapunov equation in Definition 4.4 with new terms denoted by $\Psi$ and $G$:
$BK+KB$ $=2n\Psi^{\top}\Phi K^{-1}GK^{-1}\Phi^{\top}\Psi$
$K$ $=\Psi^{\top}\Phi\Phi^{\top}\Psi$
$\Psi$ $\coloneqq\Phi K^{-\frac{1}{2}}$
$G$ $\coloneqq KC^{\top}(CKC^{\top})^{-2}CK$
Notice that $\Phi=\Psi\Psi^{\top}\Phi$. Multiplying from the left by $\Psi$ and from the right by $\Psi^{\top}$, we get for the solution $B$ of the Lyapunov equation:
$BK+KB$ $=2n\Psi^{\top}\Phi K^{-1}GK^{-1}\Phi^{\top}\Psi$
$\Psi BK\Psi^{\top}+\Psi KB\Psi^{\top}$ $=2n\Psi\Psi^{\top}\Phi K^{-1}GK^{-1}\Phi^{\top}\Psi\Psi^{\top}$
$\Psi B\Psi^{\top}\Phi\Phi^{\top}\Psi\Psi^{\top}+\Psi\Psi^{\top}\Phi\Phi^{\top}\Psi B\Psi^{\top}$ $=2n\Psi\Psi^{\top}\Phi K^{-1}GK^{-1}\Phi^{\top}\Psi\Psi^{\top}$
$\Psi B\Psi^{\top}\Phi\Phi^{\top}+\Phi\Phi^{\top}\Psi B\Psi^{\top}$ $=2n\Phi K^{-1}GK^{-1}\Phi^{\top}$ (11)
Recall Lemma C.9: See C.9
Applying this lemma to $T=\Phi K^{-\frac{1}{2}}BK^{-\frac{1}{2}}\Phi^{\top}$ (which is the form of the transformation in Definition 4.4):
$\mathcal{L}_{BT}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{\top})=0$
$\iff(2n)^{-1}A\Phi^{\top}(\Phi K^{-\frac{1}{2}}BK^{-\frac{1}{2}}\Phi^{\top}\Phi\Phi^{\top}+\Phi\Phi^{\top}\Phi K^{-\frac{1}{2}}BK^{-\frac{1}{2}}\Phi^{\top})\Phi A^{\top}=I_{d}$
$\iff(2n)^{-1}A\Phi^{\top}(\Psi B\Psi^{\top}\Phi\Phi^{\top}+\Phi\Phi^{\top}\Psi B\Psi^{\top})\Phi A^{\top}=I_{d}$
Combining with (11) we get:
$\mathcal{L}_{BT}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{\top})=0$ $\iff(2n)^{-1}A\Phi^{\top}(2n\Phi K^{-1}GK^{-1}\Phi^{\top})\Phi A^{\top}=I_{d}$
$\iff AGA^{\top}=I_{d}$
From the definition of $G$ we get that $A$ is a solution of $\mathcal{L}_{\text{BT}}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{\top})=0$ if and only if
$AKC^{\top}(CKC^{\top})^{-2}CKA^{\top}=I_{d}$ (12)
Recall Lemma C.11: See C.11
Therefore, we get the following claim:
Claim 4.
The set of solutions $A$ of $\mathcal{L}_{\text{BT}}(\{\phi_{i}\}_{i=1}^{n},\mathcal{T},A\Phi^{\top})=0$ is $\{QC:Q\in O(d)\}$.
Now we would like to use this result to show that the set of least norm minimizers of $\mathcal{L}_{\text{BT}}$ is $\{QC\Phi^{\top}:Q\in O(d)\}$. The way this follows is simple: If $f$ is a least norm minimizer of $\mathcal{L}_{\text{BT}}$, then by Proposition C.1 it has to be for the form $f=A\Phi^{\top}$ and therefore because of Claim 4 be of the form $f=QC\Phi^{\top},Q\in O(d)$. Conversely, if $f=QC\Phi^{\top},Q\in O(d)$, then Claim 4 implies it must satisfy $\mathcal{L}_{\text{BT}}(f)=0$, all it is left to prove that it is a least-norm solution, which we do next, and thereby proving the Theorem.
Assume by way of contradiction there is a $f^{\prime}\in\mathcal{F}^{\ast}$ such that $\|f^{\prime}\|_{\text{HS}}<\|f\|_{\text{HS}}$. Applying Proposition C.1 and Claim 4 we know that $f^{\prime}=\tilde{Q}^{\prime}C\Phi^{\top}$ for $\tilde{Q}^{\prime}\in O(d)$. We get that $\|\tilde{Q}C\Phi^{\top}\|_{\text{HS}}<\|\tilde{Q}^{\prime}C\Phi^{\top}\|_{\text{HS}}$ for two matrices $\tilde{Q},\tilde{Q}^{\prime}\in O(d)$. This is a contradiction since for every $Q\in O(d)$:
$\|QC\Phi^{\top}\|_{\text{HS}}^{2}=Tr((QC\Phi^{\top})^{\top}(QC\Phi^{\top}))=Tr((C\Phi^{\top})^{\top}Q^{\top}Q(C\Phi^{\top}))$
$=Tr((C\Phi^{\top})^{\top}(C\Phi^{\top})=\|C\Phi^{\top}\|_{\text{HS}}^{2}$
To summarize, we get:
$\{QC\Phi^{\top}:Q\in O(d)\}=\mathcal{F}^{\ast}$
C.4 Proof of Corollary 4.6
See 4.6
Proof.
The result is a corollary of Theorems 4.2, 4.3 and 4.5, as they imply that for a least norm minimizer of ($\mathcal{L}\Bigl{(}\{\phi(x_{i})\}_{i=1}^{n},\;\mathcal{T}_{\mathcal{H}}(FK^{-1}),\;f\Bigr{)}$ there exist an invertible matrix $M\in\mathbb{R}^{d\times d}$ and a vector $b\in\mathbb{R}^{d}$ such that:
$f=MFK^{-1}\Phi^{\top}+b$
Implying:
$f\Phi$ $=MFK^{-1}\Phi^{\top}\Phi+b$
$f\Phi$ $=MFK^{-1}K+b$
$f\Phi$ $=MF+b$
$\forall_{i\in[n]}f(\phi(x_{i}))$ $=Mf^{*}(x_{i})+b$
Appendix D The Preimage Problem for Kernel Machines
We detail closed-form approximation proposed by Honeine and Richard [2011] to the preimage problem, introduced in Section 5: Given the training data in the input space $X=[x_{1},\ldots,x_{n}]\in\mathbb{R}^{m\times n}$ and a single point $\phi^{\prime}=\Phi\theta$ in the Hilbert space, we solve the least-squares problem:
$\tilde{\phi}^{-1}(\theta)=\arg\min_{x^{\prime}}\Bigl{\|}X^{\top}x^{\prime}-(X^{\top}X-\mu_{P}K^{-1})\theta\Bigr{\|}^{2}$
where $\mu_{P}>0$ is a hyperparameter and $K$ is, as usual, the kernel Gram matrix on $X$.
There could potentially be no exact solutions $x^{\prime}$, one exact solution or multiple solutions; In our experiments, we use:
$x^{\prime}=(X^{\top})^{+}(X^{\top}X-\mu_{P}K^{-1})\theta$
Where + denotes the Moore–Penrose inverse.
Appendix E Details on the Experiments
For all of the experiments, we used either a pretrained ResNet given by ResNet50_Weights.IMAGENET1K_V2 in PyTorch [Paszke et al., 2019], or a pretrained ViT given by ViT_B_16_Weights.IMAGENET1K_V1. We reduced the dimension of the target representations to 64 using PCA.
For calculating the augmentations in Figure 2a we used 10,000 images of MNIST training data as $\{x_{i}\}_{i=1}^{n}$ and we calculated the augmented images of validation data ,which was not part of $\{x_{i}\}_{i=1}^{n}$ in Algorithm 1, meaning these are augmentations produced for new images outside of the “training set". We used $\sigma=3$ for the RBF kernel and $\lambda_{\text{ridge}}=1,\mu_{P}=1$ for these experiments ($\mu_{P}$ being a parameter for solving the preimage problem, see Section D).
For Figures 2b and 2c and the additional experiments, we used the RBF kernel with $\sigma=1$ and $\lambda=5,\mu=5,\nu=1$ for the VICReg losses. The parameters were chosen based on the empirical speed of convergence to zero Procrustes distance. We chose 10,000 images from each dataset and we optimized the loss as one batch consisting of the training images and their augmented views in the RKHS. We parameterized the learnable function as $f=C\Phi^{\top}$ where $C\in\mathbb{R}^{d\times n}$ are the learnable parameters, the calculations were done using the kernel trick. We used the Adam optimizer [Kingma and Ba, 2017] with a learning rate of 0.001. Each experiment was repeated 3 times, and the error terms correspond to the standard error of the mean.
For the original VICReg loss we used the following definition [Bardes et al., 2022]:
$L(Z,Z^{\prime})=\lambda s(Z,Z^{\prime})+\mu[v(Z)+v(Z^{\prime})]+\nu[c(Z)+c(Z^{\prime})]\quad\text{, where:}$:
$s(Z,Z^{\prime})=\frac{1}{n}\sum_{i=1}^{n}\|z_{i}-z_{i}^{\prime}\|_{2}^{2},\quad v(Z)=\frac{1}{d}\sum_{i=1}^{d}\max(0,1-\sqrt{[\text{cov}\left(Z\right)]_{i,i}+\epsilon}),$
$c(Z)=\frac{1}{d}\sum_{i\neq j}^{d}[\text{cov}(Z)]_{i,j}^{2}$
and we set $\epsilon=0.0001$.
Appendix F Extension to Neural Networks
We formalize the argument made in Section 7. Namely we show the following:
Proposition F.1 (Optimal Augmentations Guarantee Global Minimum for Neural Networks).
[…]
Proof.
We define the kernel $k_{\theta^{*}}(x,x^{\prime})=\phi_{\theta^{*}}(x)^{\top}\phi_{\theta^{*}}(x^{\prime})$. Since the rows of $\Theta^{*}$ are in the span of $\{\phi(x_{i})\}_{i=1}^{n}$, there is a matrix $C$ such that $\Theta^{*}=C[\phi_{\theta^{*}}(x_{1}),\ldots,\phi_{\theta^{*}}(x_{n})]^{T}$. The corresponding distribution of transformations $\mathcal{T}$ yields $I_{k}$ and $\Phi\,C^{\top}\bigl{(}C\,K\,C^{\top}\bigr{)}^{-1}C\,\Phi^{\top}=\Theta^{*\top}(\Theta^{*}\Theta^{*\top})^{-1}\Theta^{*}$ with probability $\frac{1}{2}$ each (Definition 4.1).
We would like to invoke the proofs of Theorems 4.2 and 4.3 with the kernel $k_{\theta^{*}}(x,x^{\prime})=\phi_{\theta^{*}}(x)^{\top}\phi_{\theta^{*}}(x^{\prime})$ and the target $C$. The proofs of the theorems imply that set of least norm minimizers of ($\mathcal{L}\Bigl{(}\{\phi_{\theta^{*}}(x_{i})\}_{i=1}^{n},\;\mathcal{T},\;\Theta\Bigr{)}$, defined as $\mathcal{F}^{\ast}$ in the proof, is not empty and its elements are of the form $A\Theta^{*}$ where $A$ is invertible. Moreover, the minimizers achieve loss 00 for VICReg and $-d$ for SCL, which are the lowest possible values the losses can take (See Remark C.7 and Lemma C.8).
However, these theorems assume that the matrix $K=[k_{\theta^{*}}(x_{i},x_{j})]_{i.j}$ is full rank (Condition 3.3). While this assumption is standard for typical kernels, in the case of neural networks it can be too strict, since usually $n>k$ (with the notable exception of the commonly studied infinitely wide limit of neural networks). Therefore, we begin by relaxing Condition 3.3 of Theorems 4.2 and 4.3.
For Theorem 4.2, Condition 3.3 is only used at the beginning of the proof of Proposition C.3 to prove that $CKC^{\top}$ is full rank. We show an alternative proof that does not assume Condition 3.3 but only uses Condition 3.2:
The matrix $K$ is psd and therefore there is a $K^{\frac{1}{2}}$ such that $K=K^{\frac{1}{2}}K^{\frac{1}{2}}$. Condition 3.2 states $\text{cov}\left([f^{\ast}(x_{1}),\ldots,f^{\ast}(x_{n})]\right)=\frac{1}{n}CKHH^{T}K^{T}C^{T}=\frac{1}{n}(CK^{\frac{1}{2}})K^{\frac{1}{2}}HH^{T}K^{T}C^{T}$ is full rank, which implies $rank(CK^{\frac{1}{2}})\geq d$ by the submultiplicativity property of the matrix rank. $C$ is a $d\times n$ matrix and therefore $rank(CK^{\frac{1}{2}})=d$. It is generally the case that the rank of a matrix $A$ is equal to the rank of the matrix $AA^{T}$ because they share the same number of non-zero singular values, therefore: $rank(CK^{\frac{1}{2}})=rank((CK^{\frac{1}{2}})(CK^{\frac{1}{2}})^{T})=rank(CKC^{T})=d$.
The proof of Theorem 4.3 relies on the proof of Theorem 4.2 to prove an equivalent result for $\mathcal{L}_{\text{VICReg-corr}}$ (Theorem C.4) and does not use the assumption that $K$ is invertible besides that. Therefore, Condition 3.3 can be relaxed for both theorems.
[…]
The Preimage Problem for Neural Networks
In the proof of Proposition F.1, we relaxed Condition 3.3 to use Theorems 4.2 and 4.3 for neural networks. However, we also implicitly assumed Condition 3.3 to calculate the preimages in the input space (Section D). Luckily, for neural networks, there is a natural way to calculate the preimages. Namely, given $\phi^{\prime}\in\mathbb{R}^{k}$ the preimage problem can be formulated as $\min_{x\in\mathcal{X}}\|\phi_{\theta^{*}}(x)-\phi^{\prime}\|^{2}$. For neural networks, the gradient $\nabla_{x}\phi_{\theta^{*}}(x)$ can be efficiently calculated. Therefore, gradient descent in the input space can be used to calculate the preimages.